### v1

# StateAct: State Tracking and Reasoning for Acting and Planning with Large Language Models

Abstract
Planning and acting to solve ‘real’ tasks using large language models (LLMs) in interactive environments has become a new frontier for AI methods. While recent advances allowed LLMs to interact with online tools, solve robotics tasks and many more, long range reasoning tasks remain a problem for LLMs. Existing methods to address this issue are very resource intensive and require additional data or human crafted rules, instead, we propose a simple method based on few-shot in-context learning alone to enhance ‘chain-of-thought’ with state-tracking for planning and acting with LLMs. We show that our method establishes the new state-of-the-art on Alfworld for in-context learning methods (+14% over the previous best few-shot in-context learning method) and performs on par with methods that use additional training data and additional tools such as code-execution. We also demonstrate that our enhanced ‘chain-of-states’ allows the agent to both solve longer horizon problems and to be more efficient in number of steps required to solve a task. We show that our method works across a variety of LLMs for both API-based and open source ones. Finally, we also conduct ablation studies and show that ‘chain-of-thoughts’ helps state-tracking accuracy, while a json-structure harms overall performance. We open-source our code and annotations at https://github.com/ai-nikolai/StateAct.
StateAct: State Tracking and Reasoning for Acting and Planning with Large Language Models
Nikolai Rozanov Imperial College London Department of Computing nikolai.rozanov13@imperial.ac.uk                        Marek Rei Imperial College London Department of Computing marek.rei@imperial.ac.uk
1 Introduction
Using the in-build world and commonsense knowledge [^1] Commonsense- and world- knowledge as explored by Lauscher et al. (2020) for example. of large language models (LLMs), such as GPT-3, Mixtral, Gemini Brown et al. (2020); Jiang et al. (2024); Anil et al. (2023) to perform interactive reasoning tasks has become a frontier in AI research, with “AI Agents” now able to solve a range of multi-modal complex tasks Durante et al. (2024). These range from solving (simulated) robotics tasks Puig et al. (2018); Shridhar et al. (2021) and digital tasks such as online shopping Yao et al. (2023a) and navigating operating systems Liu et al. (2023), to playing a variety of games Côté et al. (2019); Liu et al. (2023).
While LLMs are increasingly successful in solving AI agent tasks, Li et al. (2023b) show that LLMs struggle with long range understanding and Coelho et al. (2024) show that LLMs often focus on the beginning of the textual history, further amplifying the problem. Existing efforts to address these issues are resource intensive, Wu et al. (2024) required human expert annotations of rules, Sun et al. (2023) a code execution environment with carefully crafted code-based prompts, Fu et al. (2024) use additional training data together with retrieval augmented generation (RAG) to help the AI agent.
To overcome these challenges, we introduce “StateAct” a novel LLM agent based on few-shot in-context learning that tracks the goal by ‘reminding’ the agent of the goal and explicitly keeps track of the agent’s state (such as location and inventory). We also utilise ‘chain-of-thought’ (i.e. explicit reasoning Wei et al. (2023)) as an optional addition. At the core of the method lies a simple annotation with goal-, state- and reasoning- traces of the few-shot examples, that we call ‘chain-of-states’.
Our method establishes a new state-of-the-art for Alfworld Shridhar et al. (2021) for few-shot in-context learning based methods, outperforming the previous best method by 14% and even outperforms methods that use external tools by 2.5%. Furthermore, while our method does not require additional data or external tools and only minimal and easy human annotations, we perform on par with the current state-of-the-art that is much more resource intensive and requires additional data and expert human annotations. Comparing to in-context learning methods, we also demonstrate that our method is both more efficient in terms of number of steps to solve a task and is able to solve longer-horizon tasks, validating the hypothesis that explicit state-tracking and goal-tracking help with solving longer horizon reasoning and planning tasks.
2 Background
AI agents have historically used reinforcement learning (RL) to solve tasks Sutton and Barto (2018). With the dawn of LLMs, works such as Li et al. (2022); Nottingham et al. (2023) combined LLMs and RL and trained additional policies or value functions to make predictions.
2.1 In-context learning approaches
Huang et al. (2022a, b); Singh et al. (2022) were among the first to use LLMs directly to act in an interactive environment, their method produces agent actions as output after receiving environment observations as input.
ReAct Yao et al. (2023b) took this work further by combining ‘acting‘ Huang et al. (2022a) and ‘chain-of-thought‘ Wei et al. (2023). ReAct establishes state-of-the-art for in-context learning only based approaches and while it is a very scalable method, the performance (i.e. success rate) is still limited.
ExpeL Zhao et al. (2023) uses additional training data to generate ‘insights’ and ‘success trajectories’ during training. At inference time they look-up the closest ‘success trajectories’ as few-shot examples to the agent (as opposed to fixed few-shot examples that we use) and augment them with these ‘insights’ to perform the final inference. They achieve 59% on Alfworld using retrieved ‘success trajectories’ (as few shot examples) + ‘insights’ and 50% using the same few-shot examples as ReAct + ‘insights’.
2.1.1 In-context learning and additional tools
The current state-of-the-art for in-context learning based approaches in combination with additional tools is AdaPlanner Sun et al. (2023). They introduce a code-based prompt Li et al. (2023a) and use code-execution as an additional tool to execute the LLM generations to feed them back into the next prompt. The shortcoming of AdaPlanner is that it requires very complex human crafted prompts that are hard to scale to new environments as well as the additional step of requiring code-execution.
2.2 State tracking in LLM-based agents
Concurrent work to ours, AutoGuide Fu et al. (2024), uses ReAct as the base agent and additional training data to create ‘state-aware’ text-based guidelines for the LLM-agent, they then use a type of retrieval augmented generation (RAG) process to guide the decision making process. They embed the current observation as a summary (this is what they call ‘state-aware’) and use an LLM to ‘look up’ the relevant ‘state-aware’ guideline, which is then fed into a final LLM to generate an action. Using training data and LLM-based RAG they achieve 79.1% on top of a ReAct agent on Alfworld. Their training and RAG approach could be used complementary to our StateAct LLM agent [^2] The cost of using RAG at every generation step could be significantly higher than our method, however..
Chen et al. (2024) propose state-tracking as a way to help the agent solve the task, without training data. Their method differs to ours two-fold. Firstly, they employ a complex sequence of components working together (an LLM-based attention over the observations, an LLM-based compilation of a complex state and finally a prediction of a program). Secondly, their system involves execution of actual programs. Our method on the hand requires a straight-forward extension of ‘chain-of-thought’ and uses a single LLM call to produce the state, thought and action and we do not require program execution.
2.2.1 Fine-tuning approaches
Previous fine-tuning approaches did not significantly enhance performance Zhou et al. (2024); Yao et al. (2023b); Shridhar et al. (2021). A concurrent work, however, ActRe Yang et al. (2024b) achieves 83% by fine-tuning gpt-3.5-1106 on additional training data.
2.3 “Multi-Agent Conversation” approaches
A new trend is to use multiple LLMs concurrently to ‘chat’ to one another to produce a result. A recent work in this direction by Wu et al. (2023) achieves 67% on Alfworld.
2.4 Joined rule and LLM based agents
StateFlow Wu et al. (2024) uses Finite State Machines (FSMs) combined with LLMs to solve Alfworld. These FSMs are human-expert crafted states, transitions and rule-based heuristics, where the LLM is asked to perform limited tasks in each of the given states. While their performance is 82% on Alfworld, we believe this result is more comparable to a rule-based ‘experts’ than an LLM-based agent (notably close to 100% can be achieved on Alfworld using hand-crafted ‘experts’ Shridhar et al. (2021)).
2.5 Approaches with multiple trials
Works such as Reflexion Shinn et al. (2023) or work by Sun et al. (2023); Yang et al. (2024a) are methods that perform several evaluations on the same test environments and perform learning loops (during evaluation) on these test-set environments to improve the score. While the score necessarily goes up (as only the failed environments are re-run) we do not dive into this, as this does not compare to the actual agent performance and ‘reflection‘ can be used on top of most LLM-based agents (e.g. Reflexion uses ReAct as the base agent).
3 Method
Figure 1: StateAct interacts with the environment directly, without additional tools or resources (a). A skeleton of the StateAct prompt (b).
StateAct is an LLM-based AI agent that works on top of pre-trained large language models. StateAct takes the textual ‘observation’ from the environment and after a single call to the pre-trained LLM returns the ‘action’ back to the environment (without the use of additional tools or resources), see Figure 1(a).
StateAct utilises in-context learning Brown et al. (2020); Wei et al. (2023) to make the agent interact with the environment. At the core of the approach is a prompt that consists of few-shot examples of successful ‘interaction traces’ as well as the current ‘interaction trace’ up to the current step in the environment, see Figure 1(b). An ‘interaction trace’ consists of alternating ‘observations’ from the environment and desired (or actual) ‘outputs’ from the LLM. In the case of StateAct the LLM is tasked to generate the ‘goal’, ‘state’, ‘thought’ and ‘action’. The ‘action’ is then extracted and passed to the environment to produce the next observation, see Figure 2. For example a sample of the trace for step n in Alfworld:
Observation n:
On the countertop 1, you see a tomato 1.
LLM Output n:
>Goal: Put a clean tomato in fridge
Current Location: countertop 1
Current Inventory: None
Thought: I find the tomato, now I need to take it.
Action: take tomato 1
In the above example, the agent produces the action ‘take tomato 1’, in response to the observation ‘On the countertop 1, you see a tomato 1’.
Figure 2: An example of how StateAct is working. The depicted traces are from the Alfworld environment.
3.1 Goal-reminders and state- tracking using LLMs
In order to make StateAct more precise, we introduce simple notation. By denoting $\pi$ as the policy of an AI agent, in the standard case at time step $t$ the policy predicts action $a_{t}$, given the history of observations and actions $[o_{t},a_{t-1},...,a_{0},o_{0}]$.
$\pi(a_{t}|o_{t},a_{t-1},...,a_{0},o_{0})$ (1)
Where $a_{t}$ is the action produced by the agent at step t and $o_{t}$ is the observation produced by the environment at step t after receiving action $a_{t}$ as input. Usually, the first observation $o_{0}$ also contains the ‘goal’ description for the given environment.
For our case we need to enhance the policy to incorporate the ‘state’. Similar to previous work Yao et al. (2023b) we introduce the context vector, $c_{t}$. The context vector contains the action, as well as the other additional predictions of the agent, i.e. $c_{t}=(g_{0},s_{t},r_{t},a_{t})$. Where $g_{0}$ is the goal and always remains the same (for a given environment) and uses the goal extracted from $o_{0}$, $s_{t}$ represents the predicted state at time step $t$, $r_{t}$ represents ‘chain-of-thought’ style ‘reasoning’ at time step $t$, and $a_{t}$ represents the action at time step $t$, as before. The new policy $\pi$ then becomes:
$\pi_{contextual}(c_{t}|o_{t},c_{t-1},...,c_{0},o_{0})$ (2)
In our case the LLM acts as $\pi_{contextual}$ and produces the context vector at every time step.
4 Experimental setup
Our aim is to study long-range acting, planning and reasoning capabilities of LLM-based agents. To achieve this, in line with previous work, we turn to simulated environments as an evaluation framework and to API-based state-of-the-art large language models. Specifically, we use Alfworld Shridhar et al. (2021), a household robotics environment, and Webshop Yao et al. (2023a), an online shopping environment, as simulated environments. In line with previous work we use OpenAI’s [^3] https://openai.com, last accessed June 2024. gpt-3.5-turbo-1106. Additionally, we evaluate on additional LLMs and validate our results.
4.1 Alfworld
Figure 3: An example textual interaction in Alfworld (right) and corresponding 3D rendering (left).
Alfworld Shridhar et al. (2021) is based on a 3D, visual, household robotics environment called Alfred Shridhar et al. (2020), which was translated into a text-based environment for ease of use for language based AI models, see Figure 3. Alfworld has a total of 135 test-set examples and six environment types. It features long-time horizons, partial observability, an out-of-distribution evaluation set and text-based interactions. Alfworld simulates a household environment with a household assistant robot tasked with solving problems, e.g. clean an apple and put it on a table. The robot (or agent) then needs to perform a series of ‘high-level’ operations to accomplish the tasks, e.g. ‘go to fridge 1’, ‘open fridge 1’. At every step the environment provides a textual observation or feedback that the command has failed, e.g. ‘You open the fridge 1’, ‘You see apple 1’. The underlying text engine is based on Textworld Côté et al. (2019). See Appendix A for a complete list of commands and details on environments.
4.1.1 Alfworld correction
In our research we identified that Alfworld has a specific syntactic feature for the put command, namely put <object> in/on <place>, where “in/on” needs to be written exactly this way and using only “in” or only “on” produces a failed command. We observed this issue with LLMs on this environment and we propose a simple fix for it. We map: 1. “put <object> in <place>” and 2. “put <object> on <place>” to the command accepted by Alfworld, namely “put <object> in/on <place>”.
Methods such as AdaPlanner Sun et al. (2023) have avoided this issue because they use code-based prompts and regex parsers. However, methods such as ReAct Yao et al. (2023b) and ExpeL Zhao et al. (2023) have been affected, lowering their potential performance. In our work, we also report the results for ReAct using corrections.
4.2 Webshop
Figure 4: An example textual interaction in Webshop (right) and corresponding e-commerce website rendering (left).
Webshop Yao et al. (2023a) is a simulation of an online shopping experience. Given a task, e.g. “I want a blue water-proof winter jacket, less than \$100”, the agent needs to search a product catalogue, browse through the search results and select the most fitting product, select the attributes, e.g. colour, size, and then buy the product. In line with previous work we use the text-based version of Webshop, where all descriptions of the website are given in text form, see Figure 4. Webshop features a realistic large-scale product catalogue, a search engine and very varied product attributes depending on the category of product. See Appendix B for more details. In total the test set consists of 500 examples and each one is of the type “search and buy a product”. Overall, Webshop has a maximum of 15 steps and two commands: 1. search[<query>], 2. click[<button>].
4.3 In context learning
Since ReAct Yao et al. (2023b) forms the underlying agent for many current Zhao et al. (2023) and state-of-the-art approaches Fu et al. (2024), we use the same few-shot ‘interaction traces’ as ReAct. The main reason is to have a fair comparison and isolate additional effect - such as performance change from different in-context examples.
In total, Alfworld has six types of tasks and ReAct uses two in-context examples per task type to prompt the language models. On average each ReAct example ranges from 352 words to 591 words (590 token to 935 tokens). For our study we reuse the observations, thoughts and actions and annotate these examples further with goal ‘reminders’ and state tracking, which results in a range from 484 to 911 words (807 tokens to 1458 tokens) per example. During our annotation we discovered minor errors in the ReAct prompts and fixed them as well. We release all our annotations with our code release. In comparison, AdaPlanner Sun et al. (2023), uses a different code based approach and the prompt has 1104 words (2015 tokens) on average.
We use the two-shot examples from ReAct for Alfworld and the one-shot example from ReAct for Webshop for all our experiments.
4.4 Models
In line with previous work we focus our main experiments are based on the API based LLMs to compare performance fairly. Concretely, ‘gpt-3.5’ level models were used in previous work including current state-of-the-art ActRe (Yang et al., 2024b) and state-of-the-art for in-context learning ReAct (Yao et al., 2023b). Many OpenAI models have become deprecated. Notably, all models from ReAct and AdaPlannerSun et al. (2023) davinci-002, gpt-3.5-turbo-0301 and gpt-3.5-turbo-0613 are deprecated now. Therefore, we re-implemented ReAct and ran the experiments to determine the most suitable model, see Table 1. We establish that gpt-3.5-turbo-1106 is the best performing (from the ones that remain available) on ReAct and we therefore chose this one. In our work, we additionally compare our method using newer state-of-the-art models to show that our method generalises to new model. Specificially, we use the gpt-4o-mini and Mixtral(Jiang et al., 2024) model.. We note that full GPT-4 level models are prohibitively expensive [^4] A single evaluation run on alfworld costs approx. \$8 using gpt-3.5 and ReAct, gpt-4 would cost 10+ times more.. We use temperature 0 for all experiments and sample only the top 1 response, see Appendix C the exact settings.
Method Version AW WS
ReAct gpt-3.5-turbo-1106 40.7 18.2
ReAct gpt-3.5-turbo-0125 30.37 14.6
Table 1: ReAct success rate on Alfworld (AW) and Webshop (WS) using available gpt-3.5 models.
4.5 Metrics
In terms of metrics we use the pre-defined metrics of Alfworld and Webshop, namely success rate (SR). Success is a binary metric per each environment in the respective test sets (135 and 500 respectively). Success in Alfworld means the agent has successfully complete the whole task. In Webshop it means the agent has bought an item that has a hundred percent match with the desired item based on a partially hidden list of attributes of the shopping item (e.g. the colour, size, price, etc.).
5 Results
5.1 Alfworld
For Alfworld we present the results for ReAct, AdaPlanner with and without code execution and StateAct (ours), which consists of goal + state + thought + action. We also show StateAct without each of the components (i.e. without goal, state and thought). Interestingly we find, contrary to previous findings, that ‘thought’ or ‘reasoning’ actually sometimes harms the performance.
In Table 2, we can see that StateAct with all goal+state+thought and the correction performs the best. It outperforms ReAct with correction by around 13 points (using the same GPT model for ReAct) and by around 9 points (using the better model for ReAct). StateAct also outperforms ReAct by 22 points when corrections are not used. Furthermore, StateAct even outperforms AdaPlanner by 2.48 points, an approach that uses regex for command mapping (similar to our correction) and code-execution.
Perhaps the most surprising finding is that the simple correction described in Section 4.1.1 leads to a 16 and 23 point jump for ReAct and a 27 point jump for StateAct. This indicates that the model generally performs very well, however, struggles with minute differences in domain specific syntax.
Method GPT-3.5 SR %
Baselines w/o corrections
ReAct 0301* 51.9
ReAct 1106 40.7
ReAct 0125 30.37
ReAct (joined***) 1106 36.30
Results w/o corrections
StateAct (ours) 1106 50.37
- w/o thought 1106 62.96
- w/o goal 1106 42.96
- w/o state 1106 44.44
Baselines with corrections
ReAct + corr. 0301* 68.15
ReAct + corr. 1106 63.70
AdaPlanner No-Exec** 0301* 46.66
AdaPlanner 0301* 75.56
Results with corrections
StateAct + corr. (ours) 1106 77.04
- w/o thought 1106 64.44
- w/o goal 1106 74.04
- w/o state 1106 64.44
Table 2: Success Rate (SR) on the 135 test-set examples from Alfworld. *gpt-3.5-0301 was deprecated in June 2024. **No-Exec means AdaPlanner without code execution. ***joined means that thought + action are produced at every turn. AdaPlanner results are from Sun et al. (2023). All other results are ‘single run’.
5.2 Webshop
For Webshop we present results for ReAct and StateAct (ours). Similarly, to Alfworld we also present the results of StateAct without each of goal, state and thought. See Table 3. Interestingly, we see that removing thought produces the highest results and outperforms ReAct by 10 points. Our hypothesis is that domain specific syntax, which is more prevalent in Webshop than Alfworld, conflicts with using verbose thoughts.
Method GPT-3.5 Success Rate %
Baselines
Rule-based* N/A 9.60
RL* N/A 17.60
ReAct 1106 17.80
Results
StateAct (ours) 1106 17.00
- w/o thought 1106 27.80
- w/o goal 1106 20.40
- w/o state 1106 21.00
Table 3: Success Rate (SR) on the 500 test-set examples from Webshop. *results taken from Yao et al. (2023a). Results are ‘single run’, except for StateAct without thought and ReAct, where we ran the experiment twice and reported the average.
5.3 Results using additional models
We validate our results further by comparing on additional LLMs. Again we see a significant boost in performance, see Table 4. We ran Gpt-4o-mini and Mixtral [^5] The concrete version of Mixtral was ‘mixtral-8x22b-instruct-v0.1’. (Jiang et al., 2024) on Alfworld. ReAct achieves 68.15 and 72.59 respectively, while StateAct achieves 71.85(+3.7) and 83.70(+11.2).
Method Model SR %
Baselines
ReAct Gpt-4o-mini 68.15
ReAct Mixtral-8x22B* 72.59
Results
StateAct (ours) Gpt-4o-mini 71.85
StateAct (ours) Mixtral-8x22B* 83.70
Table 4: Success Rate (SR) on the 135 test-set examples from Alfworld. Results are ‘single run’. *=The concrete version of Mixtralis ‘mixtral-8x22b-instruct-v0.1’.
5.4 Summary of results
In conclusion we found that our simple goal-reminding and state tracking approach that purely relies on in-context learning outperforms previous in-context learning approaches by more almost 10 points and even outperforms leading approaches that rely on code-execution. Interestingly, we found that the approaches are quite sensitive to domain specific syntax and that when this is the case ‘thoughts’ that are verbose can harm performance.
6 Analysis and Ablations
In the results section we discovered that our methods perform better than previous state-of-the-art. This answers the question that we can perform better with in-context learning without resorting to additional tools, data or bigger models. In this section we want to analyse our results further and particularly also answer if our second hypothesis that goal ‘reminding’ and state tracking help with long-range reasoning actually holds. For all ablation studies we focus on Alfworld as it has two favourable properties over Webshop. Firstly, Alfworld has a longer time horizon (50 steps vs. 15 in Webshop), with tasks taking an average of less than 10 steps in Webshop and around 20 to 30 steps in Alfworld. Secondly, Alfworld has much less domain specific syntax and is purely text based, while Webshop has a more specific syntax to follow.
6.1 Do goal reminders help with long range tasks?
For this purpose we compare the original ReAct (thought+action) with just adding the goal in, i.e. StateAct (goal + thought + action). In Figure 5 we can see that while the performance of both ReAct and StateAct goes down as there are more number of steps the goal tracking has a significantly better relative performance as the number of steps increase.
To verify that this actually means that goal tracking helps with performance, as opposed to just increasing the number of steps it takes to solve a task, we calculate the average number of steps for ReAct (ignoring empty ‘thought’ turns, as otherwise ReAct would have even more steps) and StateAct. Table 5 clearly show that ReAct with an average of 38.84 steps to solve an environment is the least efficient and StateAct with an average of 28.96 steps to solve an environment is the most efficient. This shows that not only does goal tracking help with longer range tasks, it also helped with efficiency by shortening the tasks.
Figure 5: Goal vs. No Goal, on the 135 test examples from Alfworld, using gpt-3.5-turbo-1106 without correction.
6.2 What effects does state-tracking have?
We also analyse whether state tracking helps with long-range reasoning and efficiency. We compared the full StateAct against StateAct without state-tracking as well as ReAct (thought + action) against StateAct with state-tracking added (state + thought + action). In Figure 6 we see that state tracking also helps with long-range reasoning. In fact, we can see that reasoning alone is unable to solve tasks longer than 40 steps, while with state tracking even longer-range tasks can be solved than with goal-tracking alone. Also, looking at Table 5 we see that state-tracking makes the model the most efficient [^6] In terms of cost we found that despite our method using a twice longer prompt, our cost remains similar at around \$8 for the full Alfworld run, since we solve tasks more efficiently and use fewer number of steps..
Therefore we find that explicit state-tracking even further helps with long-range tasks and helps the agent solve the tasks more efficiently than without.
Figure 6: State vs. No State, on the 135 test examples from Alfworld, using gpt-3.5-turbo-1106 with correction.
Method Avg. Steps $\downarrow$
ReAct 38.84
StateAct (goal+thought+action) 31.19
StateAct (full) 28.96
ReAct with corr. 31.49
StateAct with corr. 19.11
- w/o thought 23.76
- w/o goal 20.09
Table 5: Average number of Steps (Avg. Steps) [lower is better] on the 135 test-set examples from Alfworld. gpt-3.5-1106 for all methods.
6.3 Does the model perform actual state tracking?
We ask ourselves the question if the model is actually performing state-tracking. For that purpose we look at Alfworld and construct a self-verification algorithm that is able to track the state heuristically [^7] On our few-shot prompts it achieves 100% correctness. based on the actions the agent takes. For example if the agent produces the action go to fridge 1 and the environment accepts this action we update the state with current location: fridge 1. We compare the ‘gold’ state against the predicted state. Figure 7 shows that StateAct in fact does correct state-tracking 88% of the time. We also observe that thoughts and goals help the state tracking.
Figure 7: State tracking accuracy for StateAct on 135 test examples of Alfworld using gpt-3.5-1106.
6.4 Does json structure help with performance?
Since we found that domain specific syntax harms performance, we wondered whether adding a structured format like json would help. For this purpose, we re-ran StateAct on Alfworld, but translated the state into a json format, see Appendix D for more details. Surprisingly, we found that the json format harms performance significantly, see Table 6. However, we also see that corrections help the the json format less, indicating that json helps with syntax, but harms performance.
Method SRSR (+json)%
StateAct 50.37 45.19 (-5.2)
StateAct w/ corr. 77.04 58.52(-18.5)
Table 6: Success Rate (SR) on the 135 test examples from Alfworld. Showing no-json vs. json, gpt-3.5-1106.
7 Conclusion and future work
We propose a novel method StateAct, using our ‘chain-of-states’, based on in-context learning alone and establish a new state-of-the-art for agents that do not perform training, even against methods that use code-execution. The method outperforms the previous state-of-the-art, that uses in-context learning alone, between 9% and 20% given different models and tasks and outperform in-context learning with tools (code-execution) by 3%. We also show that explicit state-tracking and goal reminders make the model more efficient as well as significantly help with longer range tasks.
We found that ‘thoughts’ or explicit reasoning do not always help performance. It would be very interesting to systematise ‘thought’ and ‘states’ and understand what contributes positively and why. Also, inspired by the good results of StateAct, it is interesting to see what other improvements can be done without resorting to training, larger model or external tools. Finally, problems related to domain specific syntax are also an interesting avenue of future work.
8 Ethical Considerations
8.1 Computational footprint
Running many of the experiments presented in this paper can have a significant computational footprint. We should consider the environment and financial resources for reproduciblity of our work. We aimed to address this concern by using gpt-3.5-turbo level models, reporting costs and minimising the cost of our method.
8.2 Hallucinations in LLMs
As LLM-based agents become more powerful and therefore more pervasive in our daily lives ‘hallucinations’ of LLMs can be very harmful Wei et al. (2024). We hope that explicit state-tracking presented in this work can also lead to future work that can reduce ‘hallucinations.’
9 Limitations
9.1 Languages and evaluation benchmarks
We evaluated our method only in the English language and on two evaluation benchmarks. While we do not expect major changes in other languages, this is something that should be investigated. Furthermore, performance on other benchmarks should be evaluated as well.
9.2 Reasoning traces rely on human judgement
Our prompts require human annotations, as such there is a natural bias present. This can have both task-performance implications as well as ethical implications.
Appendix A Alfworld
A.1 Environment Types
Alfworld has six different environment types: 1. clean, 2. heat, 3. cool, 4. examine, 5. put, 6. puttwo.
The ‘clean’ task, e.g. Task: Put a clean apple on table, requires the agent to first find the apple, then clean it (in the sinkbasin) and then put it on a table.
The ‘heat’ task, e.g. Task: Put a hot pie on table, requires the agent to first find the pie, then heat it (on the stoveburner) and then put it on a table.
The ‘cool’ task, e.g. Task: Put a cool tomato on table, requires the agent to first find the tomato, then cool it (with the fridge) and then put it on a table.
The ‘examine’ task, e.g. Task: Examine the mug with the desklamp, requires the agent to first find the mug, then find the desklamp, and then use the desklamp.
The ‘put’ task, e.g. Task: Find some apple and put it in sidetable, requires the agent to first find an apple, and then put it on the sidetable.
The ‘puttwo’ task, e.g. Task: Put two cellphone in sofa, requires the agent to first find one cellphone, and then put it on the sofa, and then to find the second one and put it on the sofa.
A.2 Action Types
Alfworld has the following valid actions: 1. go to, 2. open, 3. close, 4. put, 5. take, 6. cool, 7. heat, 8. use.
go to <place>
Example: go to table 1
open <object>
Example: open door 1
close <object>
Example: close door 1
put <object> in/on <place>
Example: put apple 1 in/on table 1
take <object> from <place>
Example: take apple 1 from table 1
cool <object> with <place>
Example: cool apple 1 with fridge 1
heat <object> with <place>
Example: heat apple 1 with fire 1
use <object>
Example: use desklamp 1
A.3 License
Alfworld has the permissible MIT license, we used it in line with the license.
Appendix B Webshop
B.1 Commands and environment
Webshop has one environment type: ‘search & buy’, as well as two commands: 1. search, 2. click.
click[<button>]
Example: click[< Back to Search]
search[<query>]
Example: search[interesting book]
B.2 Prodcuts and attributes
Webshop has over 1 million real-world products across 5 main categories (fashion, makeup, electronics, furniture, and food) and 113 sub-categories.
B.3 License
Webshop has the permissible Princeton license, we used it in line with the license.
Appendix C Code snippet to call OpenAI / GPT-3.5
client = openai.OpenAI(
# Defaults to os.environ.get("OPENAI_API_KEY")
# api_key=OPENAI_KEY,
full_prompt = [{
"role": "user",
"content": prompt
chat_completion = client.chat.completions.create(
model="gpt-3.5-turbo-1106",
messages=full_prompt,
temperature=0.0,
[…]
A prompt is given in Appendix E.
Appendix D StateAct Json Format
We translate the text based StateAct prompt:
>goal: put a hot apple in fridge
current location: starting location
current inventory: None
thought: To solve the task, I need to find and take an apple, then heat it with microwave, then put it in fridge. First I need to find an apple. An apple is more likely to appear in fridge (1), diningtable (1), coffeetable (1), drawer (1), cabinet (1-13), garbagecan (1). I can check one by one, starting with fridge 1.
action: go to fridge 1
Into the following json format:
>{"goal": "put a hot apple in fridge",
"current_location": "starting location",
"current_inventory": None,
"thought": "To solve the task, I need to find and take an apple, then heat it with microwave, then put it in fridge. First I need to find an apple. An apple is more likely to appear in fridge (1), diningtable (1), coffeetable (1), drawer (1), cabinet (1-13), garbagecan (1). I can check one by one, starting with fridge 1."
"action": "go to fridge 1"
Appendix E StateAct Prompts
E.1 Example Alfworld prompt for the ‘heat’ task
Interact with a household to solve a task.
Here are 2 examples:
You are in the middle of a room. Looking quickly around you, you see a cabinet 13, a cabinet 12, a cabinet 11, a cabinet 10, a cabinet 9, a cabinet 8, a cabinet 7, a cabinet 6, a cabinet 5, a cabinet 4, a cabinet 3, a cabinet 2, a cabinet 1, a coffeemachine 1, a countertop 1, a diningtable 1, a drawer 1, a fridge 1, a garbagecan 1, a microwave 1, a shelf 3, a shelf 2, a shelf 1, a sinkbasin 1, a stoveburner 4, a stoveburner 3, a stoveburner 2, a stoveburner 1, and a toaster 1.
Your task is to: put a hot apple in fridge.
>goal: put a hot apple in fridge
current location: starting location
current inventory: None
thought: To solve the task, I need to find and take an apple, then heat it with microwave, then put it in fridge. First I need to find an apple. An apple is more likely to appear in fridge (1), diningtable (1), coffeetable (1), drawer (1), cabinet (1-13), garbagecan (1). I can check one by one, starting with fridge 1.
action: go to fridge 1
The fridge 1 is closed.
>goal: put a hot apple in fridge
current location: fridge 1
current inventory: None
thought: None
action: open fridge 1
You open the fridge 1. The fridge 1 is open. In it, you see a cup 1, and a egg 1.
>goal: put a hot apple in fridge
current location: fridge 1
current inventory: None
thought: None
action: go to diningtable 1
On the diningtable 1, you see a apple 1, a bread 1, a fork 3, a fork 2, a fork 1, a mug 2, a peppershaker 3, a plate 2, a pot 1, a soapbottle 3, a spatula 1, a spoon 3, a spoon 2, a spoon 1, a winebottle 3, and a winebottle 2.
>goal: put a hot apple in fridge
current location: diningtable 1
current inventory: None
thought: Now I find an apple (1). Next, I need to take it.
action: take apple 1 from diningtable 1
You pick up the apple 1 from the diningtable 1.
>goal: put a hot apple in fridge
current location: diningtable 1
current inventory: apple 1
thought: Now I take an apple (1). Next, I need to go to a microwave (1) and heat it.
action: go to microwave 1
The microwave 1 is closed.
>goal: put a hot apple in fridge
current location: microwave 1
current inventory: apple 1
thought: None
action: heat apple 1 with microwave 1
You heat the apple 1 using the microwave 1.
>goal: put a hot apple in fridge
current location: microwave 1
current inventory: apple 1
thought: Now I heat an apple (1). Next, I need to put it in/on fridge 1.
action: go to fridge 1
The fridge 1 is open. In it, you see a cup 1, and a egg 1.
>goal: put a hot apple in fridge
current location: fridge 1
current inventory: apple 1
thought: None
action: put apple 1 in/on fridge 1
You are in the middle of a room. Looking quickly around you, you see a cabinet 10, a cabinet 9, a cabinet 8, a cabinet 7, a cabinet 6, a cabinet 5, a cabinet 4, a cabinet 3, a cabinet 2, a cabinet 1, a coffeemachine 1, a countertop 3, a countertop 2, a countertop 1, a diningtable 1, a drawer 6, a drawer 5, a drawer 4, a drawer 3, a drawer 2, a drawer 1, a fridge 1, a garbagecan 1, a microwave 1, a sinkbasin 1, a stoveburner 4, a stoveburner 3, a stoveburner 2, a stoveburner 1, and a toaster 1.
Your task is to: heat some egg and put it in diningtable.
>goal: heat some egg and put it in diningtable
current location: starting location
current inventory: None
thought: To solve the task, I need to find and take an egg, then heat it with microwave, then put it in diningtable. First I need to find an egg. An egg is more likely to appear in fridge (1), countertop (1-3), diningtable (1), stoveburner (1-4), toaster (1), garbagecan (1), cabinet (1-10). I can check one by one, starting with fridge 1.
action: go to fridge 1
The fridge 1 is closed.
>goal: heat some egg and put it in diningtable
current location: fridge 1
current inventory: None
thought: None
action: open fridge 1
You open the fridge 1. The fridge 1 is open. In it, you see a lettuce 2, a mug 2, and a potato 3.
>goal: heat some egg and put it in diningtable
current location: fridge 1
current inventory: None
thought: None
action: go to countertop 1
On the countertop 1, you see a bread 1, a fork 1, and a saltshaker 1.
>goal: heat some egg and put it in diningtable
current location: countertop 1
current inventory: None
thought: None
action: go to countertop 2
On the countertop 2, you see nothing.
>goal: heat some egg and put it in diningtable
current location: countertop 2
current inventory: None
thought: None
action: go to countertop 3
On the countertop 3, you see a bowl 1, a butterknife 1, a egg 2, a kettle 2, a plate 1, a sink 1, and a spatula 2.
>goal: heat some egg and put it in diningtable
current location: countertop 3
current inventory: None
thought: Now I find an egg (2). Next, I need to take it.
action: take egg 2 from countertop 3
You pick up the egg 2 from the countertop 3.
>goal: heat some egg and put it in diningtable
current location: countertop 3
current inventory: egg 2
thought: Now I take an egg (2). Next, I need go to a microwave (1) and heat it.
action: go to microwave 1
The microwave 1 is closed.
>goal: heat some egg and put it in diningtable
current location: microwave 1
current inventory: egg 2
thought: None
action: heat egg 2 with microwave 1
You heat the egg 2 using the microwave 1.
>goal: heat some egg and put it in diningtable
current location: microwave 1
current inventory: egg 2
thought: Now I heat an egg (2). Next, I need to put it in/on diningtable 1.
action: go to diningtable 1
On the diningtable 1, you see a apple 2, a bread 3, a egg 1, a kettle 1, a knife 1, a mug 1, a papertowelroll 1, a peppershaker 2, a potato 1, a soapbottle 1, and a spatula 1.
>goal: heat some egg and put it in diningtable
current location: diningtable 1
current inventory: egg 2
thought: None
action: put egg 2 in/on diningtable 1
Here is the task.
<CURRENT TASK>
E.2 Example Webshop prompt
Webshop
Instruction:
i would like a 3 ounce bottle of bright citrus deodorant for sensitive skin, and price lower than 50.00 dollars
[Search]
Goal: Buy a 3 ounce bottle of bright citrus deodorant for sensitive skin, and price lower than 50.00 dollars
Current Location: Search Home Page
Current Selection: None
Thought: None
Action: search[3 ounce bright citrus deodorant sensitive skin]
Observation:
[Back to Search]
Page 1 (Total results: 50)
[Next >]
[B078GWRC1J]
Bright Citrus Deodorant by Earth Mama | Natural and Safe for Sensitive Skin, Pregnancy and Breastfeeding, Contains Organic Calendula 3-Ounce
\$10.99
[B078GTKVXY]
Ginger Fresh Deodorant by Earth Mama | Natural and Safe for Sensitive Skin, Pregnancy and Breastfeeding, Contains Organic Calendula 3-Ounce
\$10.99
[B08KBVJ4XN]
Barrel and Oak - Aluminum-Free Deodorant, Deodorant for Men, Essential Oil-Based Scent, 24-Hour Odor Protection, Cedar & Patchouli Blend, Gentle on Sensitive Skin (Mountain Sage, 2.7 oz, 2-Pack)
\$15.95
Goal: Buy a 3 ounce bottle of bright citrus deodorant for sensitive skin, and price lower than 50.00 dollars
Current Location: Search Results Page, page 1
Current Selection: None
Thought: B078GWRC1J and B078GTKVXY are bright citrus deodorant less then 50 dollars. I can check B078GWRC1J first.
Action: click[B078GWRC1J]
Observation:
[Back to Search]
[< Prev]
scent [assorted scents][bright citrus][calming lavender][ginger fresh][simply non-scents]
size [travel set (4-pack)][3 ounce (pack of 1)][3-ounce (2-pack)]
Bright Citrus Deodorant by Earth Mama | Natural and Safe for Sensitive Skin, Pregnancy and Breastfeeding, Contains Organic Calendula 3-Ounce
Price: \$10.99
Rating: N.A.
[Description]
[Features]
[Reviews]
[Buy Now]
Goal: Buy a 3 ounce bottle of bright citrus deodorant for sensitive skin, and price lower than 50.00 dollars
Current Location: Item B078GWRC1J Page
Current Selection: None
Thought: For 3 ounce bottle of bright citrus deodorant for sensitive skin, the item has options bright citrus and ’3 ounce (pack of 1)’ and seems good to buy.
Action: click[bright citrus]
Observation: You have clicked bright citrus.
Goal: Buy a 3 ounce bottle of bright citrus deodorant for sensitive skin, and price lower than 50.00 dollars
Current Location: Item B078GWRC1J Page
Current Selection: bright citrus
Thought: None
Action: click[3 ounce (pack of 1)]
Observation: You have clicked 3 ounce (pack of 1).
Goal: Buy a 3 ounce bottle of bright citrus deodorant for sensitive skin, and price lower than 50.00 dollars
Current Location: Item B078GWRC1J Page
Current Selection: bright citrus, 3 ounce (pack of 1)
Thought: None
Action: click[Buy Now]

### v2

@@
 Planning and acting to solve ‘real’ tasks using large language models (LLMs) in interactive environments has become a new frontier for AI methods.
-While recent advances allowed LLMs to interact with online tools, solve robotics tasks and many more, long range reasoning tasks remain a problem for LLMs.
-Existing methods to address this issue are very resource intensive and require additional data or human crafted rules, instead, we propose a simple method based on few-shot in-context learning alone to enhance ‘chain-of-thought’ with state-tracking for planning and acting with LLMs.
-We show that our method establishes the new state-of-the-art on Alfworld for in-context learning methods (+14% over the previous best few-shot in-context learning method) and performs on par with methods that use additional training data and additional tools such as code-execution.
-We also demonstrate that our enhanced ‘chain-of-states’ allows the agent to both solve longer horizon problems and to be more efficient in number of steps required to solve a task.
-We show that our method works across a variety of LLMs for both API-based and open source ones.
-Finally, we also conduct ablation studies and show that ‘chain-of-thoughts’ helps state-tracking accuracy, while a json-structure harms overall performance.
+While recent advances allowed LLMs to interact with online tools and solve robotics tasks, long range reasoning tasks remain a problem for LLMs.
+Existing methods to address this issue are very resource-intensive and require additional data or human crafted rules; instead, we propose StateAct, a simple method based on few-shot in-context learning alone.
+StateAct enhances ‘chain-of-thought’ with state- and goal-tracking to help with planning and acting with LLMs.
+The method establishes the new state-of-the-art on Alfworld for in-context learning methods (+14% over the previous best few-shot in-context learning method) and performs on par with methods that use additional training data and additional tools such as code-execution.
+We also demonstrate that StateAct with enhanced ‘chain-of-states’ allows the agent to both solve longer horizon problems and to be more efficient in the number of steps required to solve a task.
+We show that StateAct works across a variety of API-based and open source LLMs.
+Finally, we also conduct ablation studies and show that ‘chain-of-thoughts’ help state-tracking accuracy, while a json-structure harms overall performance.
 We open-source our code and annotations at https://github.com/ai-nikolai/stateact.
@@
 1 Introduction
-Using the in-build world and commonsense knowledge [^1] Commonsense- and world- knowledge as explored by Lauscher et al.
-(2020) for example.
-of large language models (LLMs), such as GPT-3, Mixtral, Gemini Brown et al.
+Using the in-built world and commonsense knowledge [^1] Commonsense- and world- knowledge as explored by Lauscher et al.
+(2020), for example.
+of large language models (LLMs), such as GPT-3, Mixtral and Gemini Brown et al.
 (2020); Jiang et al.
@@
 (2024).
-These range from solving (simulated) robotics tasks Puig et al.
+These range from solving simulated robotics tasks Puig et al.
 (2018); Shridhar et al.
-(2021) and digital tasks such as online shopping Yao et al.
+(2021) and digital tasks, such as online shopping Yao et al.
 (2023a) and navigating operating systems Liu et al.
@@
 (2024) show that LLMs often focus on the beginning of the textual history, further amplifying the problem.
-Existing efforts to address these issues are resource intensive, Wu et al.
+Existing efforts to address these issues are resource-intensive: Wu et al.
 (2024) required human expert annotations of rules, Sun et al.
-(2023) a code execution environment with carefully crafted code-based prompts, Fu et al.
-(2024) use additional training data together with retrieval augmented generation (RAG) to help the AI agent.
-To overcome these challenges, we introduce “StateAct” a novel LLM agent based on few-shot in-context learning that tracks the goal by ‘reminding’ the agent of the goal and explicitly keeps track of the agent’s state (such as location and inventory).
-We also utilise ‘chain-of-thought’ (i.e.
+(2023) required a code execution environment with carefully crafted code-based prompts, and Fu et al.
+(2024) used additional training data together with retrieval augmented generation (RAG) to help the AI agent.
+To overcome these challenges, we introduce “StateAct”, a novel LLM-agent based on few-shot in-context learning.
+The method tracks the goal by ‘reminding’ the agent of the goal.
+StateAct, also, explicitly keeps track of the agent’s state (such as location and inventory).
+Additionally, we utilise ‘chain-of-thought’ (i.e.
 explicit reasoning Wei et al.
 (2023)) as an optional addition.
-At the core of the method lies a simple annotation with goal-, state- and reasoning- traces of the few-shot examples, that we call ‘chain-of-states’.
+At the core of the method lies a straightforward annotation of the few-shot examples, specifically with goal-, state- and reasoning- traces, which we call ‘chain-of-states’.
 Our method establishes a new state-of-the-art for Alfworld Shridhar et al.
-(2021) for few-shot in-context learning based methods, outperforming the previous best method by 14% and even outperforms methods that use external tools by 2.5%.
-Furthermore, while our method does not require additional data or external tools and only minimal and easy human annotations, we perform on par with the current state-of-the-art that is much more resource intensive and requires additional data and expert human annotations.
-Comparing to in-context learning methods, we also demonstrate that our method is both more efficient in terms of number of steps to solve a task and is able to solve longer-horizon tasks, validating the hypothesis that explicit state-tracking and goal-tracking help with solving longer horizon reasoning and planning tasks.
+(2021) for few-shot in-context learning-based methods, outperforming the previous best method by 14% and even outperforming methods that use external tools by 2.5%.
+Furthermore, while our method does not require additional data or external tools and only minimal and easy human annotations, we perform on par with the current state-of-the-art that is much more resource-intensive and requires additional data and expert human annotations.
+Comparing to in-context learning methods, we also demonstrate that our method is more efficient in terms of the number of steps taken to solve a task and is able to solve longer-horizon tasks, validating the hypothesis that explicit state-tracking and goal-tracking help with solving longer horizon reasoning and planning tasks.
 2 Background
 AI agents have historically used reinforcement learning (RL) to solve tasks Sutton and Barto (2018).
-With the dawn of LLMs, works such as Li et al.
-(2022); Nottingham et al.
-(2023) combined LLMs and RL and trained additional policies or value functions to make predictions.
+With the capabilities of LLMs improving, recent work has combined LLMs and RL, training additional policies or value functions to make predictions (Li et al., 2022; Nottingham et al., 2023).
 2.1 In-context learning approaches
 Huang et al.
-(2022a, b); Singh et al.
-(2022) were among the first to use LLMs directly to act in an interactive environment, their method produces agent actions as output after receiving environment observations as input.
+(2022a, b) and Singh et al.
+(2022) were among the first to use LLMs directly to act in an interactive environment; their method produces agent actions as output after receiving environment observations as input.
 ReAct Yao et al.
@@
 (2023).
-ReAct establishes state-of-the-art for in-context learning only based approaches and while it is a very scalable method, the performance (i.e.
+ReAct establishes state-of-the-art for for approaches based only on in-context learning and while it is a very scalable method, the performance (i.e.
 success rate) is still limited.
@@
 (2023) uses additional training data to generate ‘insights’ and ‘success trajectories’ during training.
-At inference time they look-up the closest ‘success trajectories’ as few-shot examples to the agent (as opposed to fixed few-shot examples that we use) and augment them with these ‘insights’ to perform the final inference.
-They achieve 59% on Alfworld using retrieved ‘success trajectories’ (as few shot examples) + ‘insights’ and 50% using the same few-shot examples as ReAct + ‘insights’.
+At inference time, they look up the closest ‘success trajectories’ as few-shot examples to the agent (as opposed to the fixed few-shot examples that we use) and augment them with these ‘insights’ to perform the final inference.
+They achieve 59% on Alfworld using retrieved ‘success trajectories’ (as few-shot examples) + ‘insights’ and 50% using the same few-shot examples as ReAct + ‘insights’.
 2.1.1 In-context learning and additional tools
-The current state-of-the-art for in-context learning based approaches in combination with additional tools is AdaPlanner Sun et al.
+The current state-of-the-art for in-context learning-based approaches in combination with additional tools is ProgPrompt (Singh et al., 2022) and follow-on work AdaPlanner Sun et al.
 (2023).
-They introduce a code-based prompt Li et al.
-(2023a) and use code-execution as an additional tool to execute the LLM generations to feed them back into the next prompt.
-The shortcoming of AdaPlanner is that it requires very complex human crafted prompts that are hard to scale to new environments as well as the additional step of requiring code-execution.
+AdaPlanner introduces a code-based prompt Li et al.
+(2023a).
+They use code-execution as an additional tool, by executing LLM generated code and feeding the results into the next LLM generation.
+The shortcoming of AdaPlanner is that it requires very complex human-crafted prompts that are hard to scale to new environments and it requires the additional step of code-execution.
 2.2 State tracking in LLM-based agents
 Concurrent work to ours, AutoGuide Fu et al.
-(2024), uses ReAct as the base agent and additional training data to create ‘state-aware’ text-based guidelines for the LLM-agent, they then use a type of retrieval augmented generation (RAG) process to guide the decision making process.
-They embed the current observation as a summary (this is what they call ‘state-aware’) and use an LLM to ‘look up’ the relevant ‘state-aware’ guideline, which is then fed into a final LLM to generate an action.
-Using training data and LLM-based RAG they achieve 79.1% on top of a ReAct agent on Alfworld.
-Their training and RAG approach could be used complementary to our StateAct LLM agent [^2] The cost of using RAG at every generation step could be significantly higher than our method, however..
+(2024), uses ReAct as the base agent as well as additional training data to create ‘state-aware’ text-based guidelines for the LLM-agent; they then use a type of retrieval augmented generation (RAG) process to guide the decision making process.
+They embed the current observation as a summary (which they call ‘state-aware’) and use an LLM to ‘look up’ the relevant ‘state-aware’ guideline, which is then fed into a final LLM to generate an action.
+Using training data and LLM-based RAG, they achieve 79.1% on top of a ReAct agent on Alfworld.
+Their training and RAG approach could be used complementary to our StateAct LLM-agent [^2] However, the cost of using RAG at every generation step could be significantly higher than our method..
 Chen et al.
-(2024) propose state-tracking as a way to help the agent solve the task, without training data.
+(2024) propose state-tracking as a way to help the agent solve the task without training data.
 Their method differs to ours two-fold.
-Firstly, they employ a complex sequence of components working together (an LLM-based attention over the observations, an LLM-based compilation of a complex state and finally a prediction of a program).
+Firstly, they employ a complex sequence of components working together (an LLM-based attention over the observations, an LLM-based compilation of a complex state and, finally, a prediction of a program).
 Secondly, their system involves execution of actual programs.
-Our method on the hand requires a straight-forward extension of ‘chain-of-thought’ and uses a single LLM call to produce the state, thought and action and we do not require program execution.
+Our method, on the other hand, requires a straightforward extension of ‘chain-of-thought’ and uses a single LLM call to produce the state, thought and action.
+Additionally, we do not require program execution.
 2.2.1 Fine-tuning approaches
@@
 A new trend is to use multiple LLMs concurrently to ‘chat’ to one another to produce a result. A recent work in this direction by Wu et al. (2023) achieves 67% on Alfworld.
-2.4 Joined rule and LLM based agents
+2.4 Joined rule and LLM-based agents
 StateFlow Wu et al.
 (2024) uses Finite State Machines (FSMs) combined with LLMs to solve Alfworld.
-These FSMs are human-expert crafted states, transitions and rule-based heuristics, where the LLM is asked to perform limited tasks in each of the given states.
-While their performance is 82% on Alfworld, we believe this result is more comparable to a rule-based ‘experts’ than an LLM-based agent (notably close to 100% can be achieved on Alfworld using hand-crafted ‘experts’ Shridhar et al.
+These FSMs are human expert-crafted states, transitions and rule-based heuristics, where the LLM is asked to perform limited tasks in each of the given states.
+While their performance is 82% on Alfworld, we believe this result is more comparable to a rule-based ‘expert’ than an LLM-based agent (notably close to 100% can be achieved on Alfworld using hand-crafted ‘experts’ Shridhar et al.
 (2021)).
@@
 (2024a) are methods that perform several evaluations on the same test environments and perform learning loops (during evaluation) on these test-set environments to improve the score.
-While the score necessarily goes up (as only the failed environments are re-run) we do not dive into this, as this does not compare to the actual agent performance and ‘reflection‘ can be used on top of most LLM-based agents (e.g.
+While the score necessarily goes up (as only the failed environments are re-run), we do not dive into this as it does not compare to the actual agent performance; additionally, ‘reflection‘ can be used on top of most LLM-based agents (e.g.
 Reflexion uses ReAct as the base agent).
@@
 StateAct is an LLM-based AI agent that works on top of pre-trained large language models.
-StateAct takes the textual ‘observation’ from the environment and after a single call to the pre-trained LLM returns the ‘action’ back to the environment (without the use of additional tools or resources), see Figure 1(a).
+StateAct takes the textual ‘observation’ from the environment and, after a single call to the pre-trained LLM, returns the ‘action’ back to the environment (without the use of additional tools or resources), see Figure 1(a).
 StateAct utilises in-context learning Brown et al.
@@
 (2023) to make the agent interact with the environment.
-At the core of the approach is a prompt that consists of few-shot examples of successful ‘interaction traces’ as well as the current ‘interaction trace’ up to the current step in the environment, see Figure 1(b).
-An ‘interaction trace’ consists of alternating ‘observations’ from the environment and desired (or actual) ‘outputs’ from the LLM.
-In the case of StateAct the LLM is tasked to generate the ‘goal’, ‘state’, ‘thought’ and ‘action’.
-The ‘action’ is then extracted and passed to the environment to produce the next observation, see Figure 2.
-For example a sample of the trace for step n in Alfworld:
+At the core of the approach is a prompt that consists of few-shot examples of successful ‘interaction traces’ as well as the current interaction trace up to the current step in the environment, see Figure 1(b).
+An interaction trace consists of alternating observations from the environment and desired (or actual) outputs from the LLM.
+In the case of StateAct, the LLM is tasked to generate the ‘goal’, ‘state’, ‘thought’ and ‘action’.
+The action is then extracted and passed to the environment to produce the next observation, see Figure 2.
+For example, a sample of the trace for step n in Alfworld is:
 Observation n:
@@
 Action: take tomato 1
-In the above example, the agent produces the action ‘take tomato 1’, in response to the observation ‘On the countertop 1, you see a tomato 1’.
+In the above example, the agent produces the action ‘take tomato 1’ in response to the observation ‘On the countertop 1, you see a tomato 1’.
 Figure 2: An example of how StateAct is working. The depicted traces are from the Alfworld environment.
 3.1 Goal-reminders and state-tracking using LLMs
-In order to make StateAct more precise, we introduce simple notation.
-By denoting $\pi$ as the policy of an AI agent, in the standard case at time step $t$ the policy predicts action $a_{t}$, given the history of observations and actions $[o_{t},a_{t-1},...,a_{0},o_{0}]$.
+In order to make StateAct more precise, we introduce goal and state tracking.
+We denote $\pi$ as the policy of an AI-agent, in the standard case at time step $t$, the policy predicts action $a_{t}$, given the history of observations and actions $[o_{t},a_{t-1},...,a_{0},o_{0}]$.
 $\pi(a_{t}|o_{t},a_{t-1},...,a_{0},o_{0})$ (1)
-Where $a_{t}$ is the action produced by the agent at step t and $o_{t}$ is the observation produced by the environment at step t after receiving action $a_{t}$ as input.
+Where $a_{t}$ is the action produced by the agent at time step t and $o_{t}$ is the observation produced by the environment at time step t after receiving action $a_{t}$ as input.
 Usually, the first observation $o_{0}$ also contains the ‘goal’ description for the given environment.
-For our case we need to enhance the policy to incorporate the ‘state’.
+For our case, we need to enhance the policy to incorporate the ‘state’.
 Similar to previous work Yao et al.
-(2023b) we introduce the context vector, $c_{t}$.
-The context vector contains the action, as well as the other additional predictions of the agent, i.e.
+(2023b), we introduce the context vector, $c_{t}$.
+The context vector contains the action, as well as the additional predictions of the agent , i.e.
 $c_{t}=(g_{0},s_{t},r_{t},a_{t})$.
-Where $g_{0}$ is the goal and always remains the same (for a given environment) and uses the goal extracted from $o_{0}$, $s_{t}$ represents the predicted state at time step $t$, $r_{t}$ represents ‘chain-of-thought’ style ‘reasoning’ at time step $t$, and $a_{t}$ represents the action at time step $t$, as before.
+Where $g_{0}$ is the goal and it always remains the same (for a given environment) and uses the goal extracted from $o_{0}$; $s_{t}$ represents the predicted state at time step $t$; $r_{t}$ represents ‘chain-of-thought’-style ‘reasoning’ at time step $t$; and $a_{t}$ represents the action at time step $t$, as before.
 The new policy $\pi$ then becomes:
 $\pi_{contextual}(c_{t}|o_{t},c_{t-1},...,c_{0},o_{0})$ (2)
-In our case the LLM acts as $\pi_{contextual}$ and produces the context vector at every time step.
+In our case, the LLM acts as $\pi_{contextual}$ and produces the context vector at every time step.
 4 Experimental setup
-Our aim is to study long-range acting, planning and reasoning capabilities of LLM-based agents.
+Our aim is to study the long-range acting, planning and reasoning capabilities of LLM-based agents.
 To achieve this, in line with previous work, we turn to simulated environments as an evaluation framework and to API-based state-of-the-art large language models.
@@
 (2021), a household robotics environment, and Webshop Yao et al.
-(2023a), an online shopping environment, as simulated environments.
-In line with previous work we use OpenAI’s [^3] https://openai.com, last accessed June 2024.
+(2023a), an online shopping environment, as our simulated environments.
+In line with previous work, we use OpenAI’s [^3] https://openai.com, last accessed June 2024.
 gpt-3.5-turbo-1106.
-Additionally, we evaluate on additional LLMs and validate our results.
+We, also, evaluate Alfworld on additional LLMs and validate our results.
 4.1 Alfworld
@@
 (2021) is based on a 3D, visual, household robotics environment called Alfred Shridhar et al.
-(2020), which was translated into a text-based environment for ease of use for language based AI models, see Figure 3.
-Alfworld has a total of 135 test-set examples and six environment types.
+(2020), which was translated into a text-based environment for ease of use for language-based AI models, see Figure 3.
+Alfworld has a total of 134 test-set examples and six environment types.
 It features long-time horizons, partial observability, an out-of-distribution evaluation set and text-based interactions.
@@
 ‘go to fridge 1’, ‘open fridge 1’.
-At every step the environment provides a textual observation or feedback that the command has failed, e.g.
+At every step, the environment provides either a textual observation or feedback that the command has failed, e.g.
 ‘You open the fridge 1’, ‘You see apple 1’.
 The underlying text engine is based on Textworld Côté et al.
-(2019).
-See Appendix A for a complete list of commands and details on environments.
+(2019), see Appendix A for a complete list of the commands and further details on the environments.
 4.1.1 Alfworld correction
-In our research we identified that Alfworld has a specific syntactic feature for the put command, namely put <object> in/on <place>, where “in/on” needs to be written exactly this way and using only “in” or only “on” produces a failed command.
-We observed this issue with LLMs on this environment and we propose a simple fix for it.
+In our research, we identified that Alfworld has a specific syntactic feature for the put command, namely put <object> in/on <place>, where “in/on” needs to be written exactly this way.
+Using only “in” or only “on” produces a failed command.
+We observed this issue with LLMs in this environment and we propose a simple fix for it.
 We map: 1.
@@
 (2023b) and ExpeL Zhao et al.
-(2023) have been affected, lowering their potential performance.
-In our work, we also report the results for ReAct using corrections.
+(2023) have been affected by it therefore lowering their potential performance.
+We report results for ReAct and StateAct both with and without these corrections.
 4.2 Webshop
-Figure 4: An example textual interaction in Webshop (right) and corresponding e-commerce website rendering (left).
+Figure 4: An textual interaction in Webshop (right) and correspondingwebsite rendering (left).
 Webshop Yao et al.
@@
 Given a task, e.g.
-“I want a blue water-proof winter jacket, less than \$100”, the agent needs to search a product catalogue, browse through the search results and select the most fitting product, select the attributes, e.g.
+“I want a blue water-proof winter jacket, less than \$100”, the agent needs to search a product catalogue, browse through the search results, select the most fitting product, select the attributes, e.g.
 colour, size, and then buy the product.
-In line with previous work we use the text-based version of Webshop, where all descriptions of the website are given in text form, see Figure 4.
-Webshop features a realistic large-scale product catalogue, a search engine and very varied product attributes depending on the category of product.
-See Appendix B for more details.
-In total the test set consists of 500 examples and each one is of the type “search and buy a product”.
+In line with previous work, we use the text-based version of Webshop, where all descriptions of the website are given in text form, see Figure 4.
+Webshop features a realistic large-scale product catalogue, a search engine and very varied product attributes depending on the category of product, see Appendix B for more details.
+In total, the test set consists of 500 examples and each one is of the type “search and buy a product”.
 Overall, Webshop has a maximum of 15 steps and two commands: 1.
@@
 (2024), we use the same few-shot ‘interaction traces’ as ReAct.
-The main reason is to have a fair comparison and isolate additional effect - such as performance change from different in-context examples.
+The main reason is to have a fair comparison and to isolate additional effects - such as performance change - from different in-context examples.
 In total, Alfworld has six types of tasks and ReAct uses two in-context examples per task type to prompt the language models.
-On average each ReAct example ranges from 352 words to 591 words (590 token to 935 tokens).
-For our study we reuse the observations, thoughts and actions and annotate these examples further with goal ‘reminders’ and state tracking, which results in a range from 484 to 911 words (807 tokens to 1458 tokens) per example.
-During our annotation we discovered minor errors in the ReAct prompts and fixed them as well.
+On average, each ReAct example ranges from 352 words to 591 words (590 tokens to 935 tokens).
+For our study, we reuse the observations, thoughts and actions, and annotate these examples further with goal ‘reminders’ and state tracking, which results in a range from 484 to 911 words (807 tokens to 1458 tokens) per example.
+During our annotation, we discovered minor errors in the ReAct prompts and fixed them as well.
 We release all our annotations with our code release.
 In comparison, AdaPlanner Sun et al.
-(2023), uses a different code based approach and the prompt has 1104 words (2015 tokens) on average.
-We use the two-shot examples from ReAct for Alfworld and the one-shot example from ReAct for Webshop for all our experiments.
+(2023), uses a different code-based approach and the prompt has 1104 words (2015 tokens) on average.
+We use the two-shot examples from ReAct for Alfworld and the one-shot example from ReAct for Webshop in all our experiments.
 4.4 Models
-In line with previous work we focus our main experiments are based on the API based LLMs to compare performance fairly.
+In line with previous work, we focus our main experiments are based on the API-based LLMs to compare performance fairly.
 Concretely, ‘gpt-3.5’ level models were used in previous work including current state-of-the-art ActRe (Yang et al., 2024b) and state-of-the-art for in-context learning ReAct (Yao et al., 2023b).
-Many OpenAI models have become deprecated.
-Notably, all models from ReAct and AdaPlannerSun et al.
-(2023) davinci-002, gpt-3.5-turbo-0301 and gpt-3.5-turbo-0613 are deprecated now.
+Many OpenAI models have become deprecated; notably, all models from ReAct and AdaPlanner Sun et al.
+(2023) davinci-002, gpt-3.5-turbo-0301, and gpt-3.5-turbo-0613 are now deprecated.
 Therefore, we re-implemented ReAct and ran the experiments to determine the most suitable model, see Table 1.
-We establish that gpt-3.5-turbo-1106 is the best performing (from the ones that remain available) on ReAct and we therefore chose this one.
-In our work, we additionally compare our method using newer state-of-the-art models to show that our method generalises to new model.
-Specificially, we use the gpt-4o-mini and Mixtral(Jiang et al., 2024) model..
+We establish that gpt-3.5-turbo-1106 is the best performing (from the ones that remain available) on ReAct and we therefore chose this model for our main experiments.
+In our work, we also compare our method using newer state-of-the-art models to show that our method generalises to new models.
+Specificially, we use the gpt-4o-mini [^4] Again, we chose the smaller model due to cost.
+and Mixtral (Jiang et al., 2024) models.
 We note that full GPT-4 level models are prohibitively expensive [^5] A single evaluation run on Alfworld costs approx.
-\$8 using gpt-3.5 and ReAct, gpt-4 would cost 10+ times more..
-We use temperature 0 for all experiments and sample only the top 1 response, see Appendix C the exact settings.
+\$8 using gpt-3.5 and ReAct, whilst gpt-4 costs 10+ times more..
+We use temperature 0 for all experiments and sample only the top 1 response, see Appendix C for the exact settings.
 Method Version AW WS
@@
 ReAct gpt-3.5-turbo-0125 30.37 14.6
-Table 1: ReAct success rate on Alfworld (AW) and Webshop (WS) using available gpt-3.5 models.
+Table 1: ReAct success rate on Alfworld (AW) and Webshop (WS), using available gpt-3.5 models.
 4.5 Metrics
-In terms of metrics we use the pre-defined metrics of Alfworld and Webshop, namely success rate (SR).
-Success is a binary metric per each environment in the respective test sets (135 and 500 respectively).
-Success in Alfworld means the agent has successfully complete the whole task.
-In Webshop it means the agent has bought an item that has a hundred percent match with the desired item based on a partially hidden list of attributes of the shopping item (e.g.
+In terms of metrics, we use the pre-defined metrics of Alfworld and Webshop, namely success rate (SR).
+Success is a binary metric per each environment in the respective test sets (134 for Alfworld and 500 for Webshop respectively).
+Success in Alfworld means the agent has successfully completed the whole task.
+In Webshop, it means the agent has bought an item that has a one hundred percent match with the desired item based on a partially hidden list of shopping item attributes (e.g.
 the colour, size, price, etc.).
@@
 5.1 Alfworld
-For Alfworld we present the results for ReAct, AdaPlanner with and without code execution and StateAct (ours), which consists of goal + state + thought + action.
+For Alfworld, we present the results for ReAct, AdaPlanner with and without code execution, and StateAct (ours), which consists of goal + state + thought + action.
 We also show StateAct without each of the components (i.e.
 without goal, state and thought).
-Interestingly we find, contrary to previous findings, that ‘thought’ or ‘reasoning’ actually sometimes harms the performance.
+Interestingly, we find that, contrary to previous findings, ‘thought’ or ‘reasoning’ actually sometimes hinders the performance.
 In Table 2, we can see that StateAct with all goal + state + thought and the correction performs the best.
-It outperforms ReAct with correction by around 13 points (using the same GPT model for ReAct) and by around 9 points (using the better model for ReAct).
+It outperforms ReAct with correction by around 13 points (when using the same GPT model for ReAct) and by around 9 points (when using the better model for ReAct).
 StateAct also outperforms ReAct by 22 points when corrections are not used.
-Furthermore, StateAct even outperforms AdaPlanner by 2.48 points, an approach that uses regex for command mapping (similar to our correction) and code-execution.
-Perhaps the most surprising finding is that the simple correction described in Section 4.1.1 leads to a 16 and 23 point jump for ReAct and a 27 point jump for StateAct.
-This indicates that the model generally performs very well, however, struggles with minute differences in domain specific syntax.
+Furthermore, StateAct even outperforms AdaPlanner by 2.48 points, an approach that uses regex for command-mapping (similar to our correction) along with using code-execution.
+Perhaps the most surprising finding is that the simple correction described in Section 4.1.1 leads to a 16- and 23-point jump for ReAct and a 27-point jump for StateAct.
+This indicates that the model generally performs very well; however, it struggles with minute differences in domain-specific syntax.
 Method GPT-3.5 SR %
@@
 - w/o state 1106 64.44
-Table 2: Success Rate (SR) on the 135 test-set examples from Alfworld.
+Table 2: Success Rate (SR) on the 134 test-set examples from Alfworld.
 *gpt-3.5-0301 was deprecated in June 2024.
@@
 (2023).
-All other results are ‘single run’.
+All other results are single run.
 5.2 Webshop
-For Webshop we present results for ReAct and StateAct (ours).
-Similarly, to Alfworld we also present the results of StateAct without each of goal, state and thought.
-See Table 3.
+For Webshop, we present the results for ReAct and StateAct.
+Similarly to Alfworld, we also present the results of StateAct without each of goal, state and thought, see Table 3.
 Interestingly, we see that removing thought produces the highest results and outperforms ReAct by 10 points.
-Our hypothesis is that domain specific syntax, which is more prevalent in Webshop than Alfworld, conflicts with using verbose thoughts.
+Our hypothesis is that domain-specific syntax, which is more prevalent in Webshop than Alfworld, conflicts with using verbose thoughts.
 Method GPT-3.5 Success Rate %
@@
 (2023a).
-Results are ‘single run’, except for StateAct without thought and ReAct, where we ran the experiment twice and reported the average.
+For StateAct without thought and ReAct Results are the average of two runs.
+Other Results are single run.
 5.3 Results using additional models
-We validate our results further by comparing on additional LLMs.
+We validate our results further with additional LLMs.
 Again we see a significant boost in performance, see Table 4.
-We ran Gpt-4o-mini and Mixtral [^5] The concrete version of Mixtral was ‘mixtral-8x22b-instruct-v0.1’.
+We experimented with gpt-4o-mini and Mixtral [^6] The concrete version of Mixtral was ‘mixtral-8x22b-instruct-v0.1’.
 (Jiang et al., 2024) on Alfworld.
@@
 StateAct (ours) Mixtral-8x22B* 83.70
-Table 4: Success Rate (SR) on the 135 test-set examples from Alfworld. Results are ‘single run’. *=The concrete version of Mixtralis ‘mixtral-8x22b-instruct-v0.1’.
+Table 4: Success Rate (SR) on the 134 test-set examples from Alfworld. Results are ‘single run’. *=The concrete version of Mixtral is ‘mixtral-8x22b-instruct-v0.1’.
 5.4 Summary of results
-In conclusion we found that our simple goal-reminding and state tracking approach that purely relies on in-context learning outperforms previous in-context learning approaches by more almost 10 points and even outperforms leading approaches that rely on code-execution.
-Interestingly, we found that the approaches are quite sensitive to domain specific syntax and that when this is the case ‘thoughts’ that are verbose can harm performance.
+In conclusion, we found that our simple goal-reminding and state-tracking approach that purely relies on in-context learning outperforms previous in-context learning approaches by more than 10 points and even outperforms leading approaches that rely on code-execution.
+Interestingly, we found that the approaches are quite sensitive to domain-specific syntax and, when this is the case, ‘thoughts’ that are verbose can hinder performance.
 6 Analysis and Ablations
-In the results section we discovered that our methods perform better than previous state-of-the-art.
-This answers the question that we can perform better with in-context learning without resorting to additional tools, data or bigger models.
-In this section we want to analyse our results further and particularly also answer if our second hypothesis that goal ‘reminding’ and state tracking help with long-range reasoning actually holds.
-For all ablation studies we focus on Alfworld as it has two favourable properties over Webshop.
+In the results section, we discovered that our methods perform better than previous state-of-the-art.
+Our method shows strong performance with in-context learning without resorting to additional tools, data or bigger models.
+In this section, we analyse our results further and also answer our second hypothesis that goal ‘reminding’ and state-tracking help with long-range reasoning.
+For all ablation studies, we focus on Alfworld as it has two favourable properties over Webshop.
 Firstly, Alfworld has a longer time horizon (50 steps vs.
-15 in Webshop), with tasks taking an average of less than 10 steps in Webshop and around 20 to 30 steps in Alfworld.
-Secondly, Alfworld has much less domain specific syntax and is purely text based, while Webshop has a more specific syntax to follow.
-6.1 Do goal reminders help with long range tasks?
-For this purpose we compare the original ReAct (thought+action) with just adding the goal in, i.e.
+15 in Webshop), with tasks taking, on average, less than 10 steps in Webshop and around 20 to 30 steps in Alfworld.
+Secondly, Alfworld has much less domain-specific syntax and is purely text-based, while Webshop has a more specific syntax to follow.
+6.1 Do goal reminders help with long-range tasks?
+For this purpose, we compare the original ReAct (thought + action) with the goal included, i.e.
 StateAct (goal + thought + action).
-In Figure 5 we can see that while the performance of both ReAct and StateAct goes down as there are more number of steps the goal tracking has a significantly better relative performance as the number of steps increase.
-To verify that this actually means that goal tracking helps with performance, as opposed to just increasing the number of steps it takes to solve a task, we calculate the average number of steps for ReAct (ignoring empty ‘thought’ turns, as otherwise ReAct would have even more steps) and StateAct.
-Table 5 clearly show that ReAct with an average of 38.84 steps to solve an environment is the least efficient and StateAct with an average of 28.96 steps to solve an environment is the most efficient.
-This shows that not only does goal tracking help with longer range tasks, it also helped with efficiency by shortening the tasks.
-Figure 5: Goal vs. No Goal, on the 135 test examples from Alfworld, using gpt-3.5-turbo-1106 without correction.
+In Figure 5 we can see that, while the performance of both ReAct and StateAct goes down as there are more steps, the goal tracking has a substantially better relative performance as the number of steps increases.
+This finding is in-line with our original motivation that LLM agents deteriorate in performance as the prompts and interactions get longer.
+To verify that this actually means that goal tracking helps with performance, as opposed to just increasing the number of steps it takes to solve a task, we calculate the average number of steps for ReAct [^7] We ignore ‘thought’ turns for ReAct as otherwise ReAct would have even more steps.
+and StateAct.
+Table 5 clearly show that ReAct, with an average of 38.84 steps to solve an environment, is the least efficient whilst StateAct, with an average of 28.96 steps to solve an environment, is the most efficient.
+This shows that not only does goal reminders help with longer range tasks, it also helps with efficiency, by shortening the tasks.
+See Appendix E for more discussion.
+Figure 5: Goal vs. No Goal, on the test set of Alfworld, using gpt-3.5-turbo-1106 without correction.
 6.2 What effects does state-tracking have?
-We also analyse whether state tracking helps with long-range reasoning and efficiency.
-We compared the full StateAct against StateAct without state-tracking as well as ReAct (thought + action) against StateAct with state-tracking added (state + thought + action).
-In Figure 6 we see that state tracking also helps with long-range reasoning.
-In fact, we can see that reasoning alone is unable to solve tasks longer than 40 steps, while with state tracking even longer-range tasks can be solved than with goal-tracking alone.
-Also, looking at Table 5 we see that state-tracking makes the model the most efficient [^6] In terms of cost we found that despite our method using a twice longer prompt, our cost remains similar at around \$8 for the full Alfworld run, since we solve tasks more efficiently and use fewer number of steps..
-Therefore we find that explicit state-tracking even further helps with long-range tasks and helps the agent solve the tasks more efficiently than without.
-Figure 6: State vs. No State, on the 135 test examples from Alfworld, using gpt-3.5-turbo-1106 with correction.
+We also analyse whether state-tracking helps with long-range reasoning and efficiency.
+We compare the full StateAct against StateAct without state-tracking, as well as comparing ReAct (thought + action) against StateAct with state-tracking (state + thought + action).
+In Figure 6 we see that state-tracking also helps with long-range reasoning.
+In fact, we can see that reasoning alone is unable to solve tasks longer than 40 steps, while with both state tracking and goal-tracking longer-range tasks can be solved.
+Also, looking at Table 5 we see that state-tracking makes the model the most efficient [^8] In terms of cost, we found that, despite our method using a twice-longer prompt, our cost remains similar, at around \$8 for the full Alfworld run, since we solve tasks more efficiently and use fewer steps..
+Concretely, ReAct has 24/134 examples that are in the bucket ‘40-50’ and solves 0, while our method has 6/134 examples and solves 4.
+Therefore, we find that explicit state-tracking helps with long-range tasks and enables the agent to solve the tasks more efficiently.
+Figure 6: State vs. No State, on the 134 test examples from Alfworld. gpt-3.5-turbo-1106 with correction.
 Method Avg. Steps $\downarrow$
@@
 - w/o goal 20.09
-Table 5: Average number of Steps (Avg. Steps) [lower is better] on the 135 test-set examples from Alfworld. gpt-3.5-1106 for all methods.
-6.3 Does the model perform actual state tracking?
-We ask ourselves the question if the model is actually performing state-tracking.
-For that purpose we look at Alfworld and construct a self-verification algorithm that is able to track the state heuristically [^7] On our few-shot prompts it achieves 100% correctness.
+Table 5: Average number of Steps (Avg. Steps) [lower is better] on test-set of Alfworld, using gpt-3.5-1106.
+6.3 Does the model perform actual state-tracking?
+We ask ourselves whether the model is actually performing state-tracking.
+For that purpose, we look at Alfworld and construct a self-verification algorithm that is able to track the state heuristically [^9] See Appendix D for details.
 based on the actions the agent takes.
-For example if the agent produces the action go to fridge 1 and the environment accepts this action we update the state with current location: fridge 1.
+For example, if the agent produces the action go to fridge 1 and the environment accepts this action, we update the state with current location: fridge 1.
 We compare the ‘gold’ state against the predicted state.
-Figure 7 shows that StateAct in fact does correct state-tracking 88% of the time.
-We also observe that thoughts and goals help the state tracking.
-Figure 7: State tracking accuracy for StateAct on 135 test examples of Alfworld using gpt-3.5-1106.
+Figure 7 shows that StateAct does in fact correct state-tracking 88% of the time.
+We also observe that thoughts and goals help the state-tracking.
+Figure 7: State-tracking accuracy for StateAct on 134 test examples of Alfworld, using gpt-3.5-1106.
 6.4 Does json structure help with performance?
-Since we found that domain specific syntax harms performance, we wondered whether adding a structured format like json would help.
-For this purpose, we re-ran StateAct on Alfworld, but translated the state into a json format, see Appendix D for more details.
-Surprisingly, we found that the json format harms performance significantly, see Table 6.
-However, we also see that corrections help the the json format less, indicating that json helps with syntax, but harms performance.
+Since we found that domain-specific syntax hinders performance, we investigated whether adding a structured format like json would help.
+For this purpose, we re-ran StateAct on Alfworld, but translated the state into a json format, see Appendix G for more details.
+Surprisingly, we found that the json format hinders performance significantly, see Table 6.
+However, we also see that corrections help the json format less, indicating that json helps with syntax but hinders performance.
 Method SRSR (+json)%
@@
 StateAct w/ corr. 77.04 58.52(-18.5)
-Table 6: Success Rate (SR) on the 135 test examples from Alfworld. Showing no-json vs. json, gpt-3.5-1106.
+Table 6: No-json vs. json. Success Rate (SR) on the test set of Alfworld, using gpt-3.5-1106.
 7 Conclusion and future work
-We propose a novel method StateAct, using our ‘chain-of-states’, based on in-context learning alone and establish a new state-of-the-art for agents that do not perform training, even against methods that use code-execution.
-The method outperforms the previous state-of-the-art, that uses in-context learning alone, between 9% and 20% given different models and tasks and outperform in-context learning with tools (code-execution) by 3%.
-We also show that explicit state-tracking and goal reminders make the model more efficient as well as significantly help with longer range tasks.
-We found that ‘thoughts’ or explicit reasoning do not always help performance.
-It would be very interesting to systematise ‘thought’ and ‘states’ and understand what contributes positively and why.
-Also, inspired by the good results of StateAct, it is interesting to see what other improvements can be done without resorting to training, larger model or external tools.
-Finally, problems related to domain specific syntax are also an interesting avenue of future work.
+Inspired by the challenge that LLMs and LLM agents struggle with long context, we propose a novel in context learning method for LLM agents - StateAct.
+We introduce ‘chain-of-states’, which relies on in-context learning alone and establish a new state-of-the-art for agents that do not perform training, even against methods that use code-execution.
+StateAct outperforms the previous state-of-the-art that uses in-context learning alone, by between 9% and 20% given different models and tasks, and outperforms in-context learning with tools (i.e.
+code-execution) by 3%, while performing on par with methods that require additional training data.
+In conclusion, we show that explicit state-tracking and goal reminders help the model solve long range reasoning tasks and interestingly also make the model more efficient.
 8 Ethical Considerations
@@
 8.2 Hallucinations in LLMs
-As LLM-based agents become more powerful and therefore more pervasive in our daily lives ‘hallucinations’ of LLMs can be very harmful Wei et al.
+As LLM-based agents become more powerful and therefore more pervasive in our daily lives, ‘hallucinations’ in LLMs can be very harmful Wei et al.
 (2024).
-We hope that explicit state-tracking presented in this work can also lead to future work that can reduce ‘hallucinations.’
+We hope that explicit state-tracking presented in this work can also lead to future work to reduce ‘hallucinations.’
 9 Limitations
@@
 Furthermore, performance on other benchmarks should be evaluated as well.
-9.2 Reasoning traces rely on human judgement
-Our prompts require human annotations, as such there is a natural bias present. This can have both task-performance implications as well as ethical implications.
+9.2 Reasoning traces that rely on human judgement
+Our prompts require human annotations; as such, there is a natural bias present. This can have both task-performance implications as well as ethical implications.
 Appendix A Alfworld
@@
 Alfworld has six different environment types: 1. clean, 2. heat, 3. cool, 4. examine, 5. put, 6. puttwo.
-The ‘clean’ task, e.g. Task: Put a clean apple on table, requires the agent to first find the apple, then clean it (in the sinkbasin) and then put it on a table.
-The ‘heat’ task, e.g. Task: Put a hot pie on table, requires the agent to first find the pie, then heat it (on the stoveburner) and then put it on a table.
+The ‘clean’ task, e.g. Task: Put a clean apple on table, requires the agent to first find the apple, then clean it (in the sink/basin) and then put it on a table.
+The ‘heat’ task, e.g. Task: Put a hot pie on table, requires the agent to first find the pie, then heat it (on the stove/burner) and then put it on a table.
 The ‘cool’ task, e.g. Task: Put a cool tomato on table, requires the agent to first find the tomato, then cool it (with the fridge) and then put it on a table.
@@
 A.3 License
-Alfworld has the permissible MIT license, we used it in line with the license.
+Alfworld has the permissible MIT license; we used it in line with the license.
 Appendix B Webshop
@@
 B.2 Prodcuts and attributes
-Webshop has over 1 million real-world products across 5 main categories (fashion, makeup, electronics, furniture, and food) and 113 sub-categories.
+Webshop has over 1 million real-world products across 5 main categories (fashion, makeup, electronics, furniture and food) and 113 sub-categories.
 B.3 License
-Webshop has the permissible Princeton license, we used it in line with the license.
+Webshop has the permissible Princeton license; we used it in line with the license.
 Appendix C Code snippet to call OpenAI / GPT-3.5
@@
 […]
-A prompt is given in Appendix E.
-Appendix D StateAct Json Format
-We translate the text based StateAct prompt:
+A prompt is given in Appendix H.
+Appendix D Heuristic State-tracking Explained
+Heuristic state tracking is based on the idea that the state can be inferred automatically if one follows the actions of the agent and observations of the environments.
+Specifically, the state at time $t$ for StateAct depends on the state at time t-1 and the action $a_{t-1}$ and observation $o_{t}$.
+For example, if the state at time $t$ the ‘current location’ of the state is set to table 1 and the action is go to fridge 1 and the observation is successful, then the ‘current location’ can be updated to be fridge 1 automatically.
+This rule based ‘state-tracking’ is how the heuristics work.
+Appendix E ‘Step Length Analysis’ Discussion
+An alternative to calculating and comparing step length could be ‘gold solutions’ to measure optimal step length and optimality of an agent.
+We see two issues.
+Firstly, the annotation cost of creating gold solutions.
+Secondly, it is not clear what the gold solution should be.
+Concretely in Alfworld, an ‘oracle’ solution could have very few steps as it would immediately go to the location of the ‘hidden’ object, while a ‘non-oracle’ expert solution would have more steps as more locations would be searched.
+Thus a ‘difficulty’ measure would be needed instead, but it is ambiguous.
+Appendix F Potential Future Work Directions
+We found that ‘thoughts’ or explicit reasoning do not always help performance.
+It would be very interesting to systematise ‘thought’ and ‘states’ and to understand what contributes positively and the reasons why.
+Also, inspired by the positive results of StateAct, it is interesting to see what other improvements can be made without resorting to training, larger models or external tools.
+Finally, problems related to domain-specific syntax are also an interesting avenue for future work.
+Appendix G StateAct Json Format
+We translate the text-based StateAct prompt:
 >goal: put a hot apple in fridge
@@
 "action": "go to fridge 1"
-Appendix E StateAct Prompts
-E.1 Example Alfworld prompt for the ‘heat’ task
+Appendix H StateAct Prompts
+H.1 Example Alfworld prompt for the ‘heat’ task
 Interact with a household to solve a task.
@@
 <CURRENT TASK>
-E.2 Example Webshop prompt
+H.2 Example Webshop prompt
 Webshop

### v3

@@
-StateAct: State Tracking and Reasoning for Acting and Planning with Large Language Models
+StateAct: Enhancing LLM Base Agents via Self-prompting and State-tracking
 Abstract
-Planning and acting to solve ‘real’ tasks using large language models (LLMs) in interactive environments has become a new frontier for AI methods.
-While recent advances allowed LLMs to interact with online tools and solve robotics tasks, long range reasoning tasks remain a problem for LLMs.
-Existing methods to address this issue are very resource-intensive and require additional data or human crafted rules; instead, we propose StateAct, a simple method based on few-shot in-context learning alone.
-StateAct enhances ‘chain-of-thought’ with state- and goal-tracking to help with planning and acting with LLMs.
-The method establishes the new state-of-the-art on Alfworld for in-context learning methods (+14% over the previous best few-shot in-context learning method) and performs on par with methods that use additional training data and additional tools such as code-execution.
-We also demonstrate that StateAct with enhanced ‘chain-of-states’ allows the agent to both solve longer horizon problems and to be more efficient in the number of steps required to solve a task.
-We show that StateAct works across a variety of API-based and open source LLMs.
-Finally, we also conduct ablation studies and show that ‘chain-of-thoughts’ help state-tracking accuracy, while a json-structure harms overall performance.
-We open-source our code and annotations at https://github.com/ai-nikolai/stateact.
-StateAct: State Tracking and Reasoning for Acting and Planning with Large Language Models
+Large language models (LLMs) are increasingly used as autonomous agents, tackling tasks from robotics to web navigation.
+Their performance depends on the underlying base agent.
+Existing methods, however, struggle with long-context reasoning and goal adherence.
+We introduce StateAct, a novel and efficient base agent that enhances decision-making through (1) self-prompting, which reinforces task goals at every step, and (2) chain-of-states, an extension of chain-of-thought that tracks state information over time.
+StateAct outperforms ReAct, the previous best base agent, by over 10% on Alfworld, 30% on Textcraft, and 7% on Webshop across multiple frontier LLMs.
+We also demonstrate that StateAct can be used as a drop-in replacement for ReAct with advanced LLM agent methods such as test-time scaling, yielding an additional 12% gain on Textcraft.
+By improving efficiency and long-range reasoning without requiring additional training or retrieval, StateAct provides a scalable foundation for LLM agents.
+We open source our code to support further research at https://github.com/ai-nikolai/stateact.
+StateAct: Enhancing LLM Base Agents via Self-prompting and State-tracking
 Nikolai Rozanov Imperial College London Department of Computing nikolai.rozanov13@imperial.ac.uk                        Marek Rei Imperial College London Department of Computing marek.rei@imperial.ac.uk
 1 Introduction
-Using the in-built world and commonsense knowledge [^1] Commonsense- and world- knowledge as explored by Lauscher et al.
+Leveraging the in-built world and commonsense knowledge [^1] Commonsense- and world- knowledge as explored by Lauscher et al.
 (2020), for example.
-of large language models (LLMs), such as GPT-3, Mixtral and Gemini Brown et al.
+of large language models (LLMs), such as GPT, Gemini and Mixtral Brown et al.
+(2020); Anil et al.
 (2023); Jiang et al.
-(2024); Anil et al.
-(2023) to perform interactive reasoning tasks has become a frontier in AI research, with “AI Agents” now able to solve a range of multi-modal complex tasks Durante et al.
+(2024) to perform interactive reasoning tasks has become a frontier in AI research.
+“AI Agents” are now able to solve a range of multi-modal complex tasks Durante et al.
 (2024).
-These range from solving simulated robotics tasks Puig et al.
+These include simulated robotics tasks Puig et al.
 (2018); Shridhar et al.
 (2021) and digital tasks, such as online shopping Yao et al.
-(2023a) and navigating operating systems Liu et al.
-(2023), to playing a variety of games Côté et al.
+(2023a), navigating operating systems Liu et al.
+(2023), and playing a variety of games Côté et al.
 (2019); Liu et al.
+(2023); Prasad et al.
 (2024).
-While LLMs are increasingly successful in solving AI agent tasks, Li et al.
-(2023b) show that LLMs struggle with long range understanding and Coelho et al.
-(2024) show that LLMs often focus on the beginning of the textual history, further amplifying the problem.
-Existing efforts to address these issues are resource-intensive: Wu et al.
-(2024) required human expert annotations of rules, Sun et al.
-(2023) required a code execution environment with carefully crafted code-based prompts, and Fu et al.
-(2024) used additional training data together with retrieval augmented generation (RAG) to help the AI agent.
-To overcome these challenges, we introduce “StateAct”, a novel LLM-agent based on few-shot in-context learning.
-The method tracks the goal by ‘reminding’ the agent of the goal.
-StateAct, also, explicitly keeps track of the agent’s state (such as location and inventory).
-Additionally, we utilise ‘chain-of-thought’ (i.e.
-explicit reasoning Wei et al.
-(2023)) as an optional addition.
-At the core of the method lies a straightforward annotation of the few-shot examples, specifically with goal-, state- and reasoning- traces, which we call ‘chain-of-states’.
-Our method establishes a new state-of-the-art for Alfworld Shridhar et al.
-(2021) for few-shot in-context learning-based methods, outperforming the previous best method by 14% and even outperforming methods that use external tools by 2.5%.
-Furthermore, while our method does not require additional data or external tools and only minimal and easy human annotations, we perform on par with the current state-of-the-art that is much more resource-intensive and requires additional data and expert human annotations.
-Comparing to in-context learning methods, we also demonstrate that our method is more efficient in terms of the number of steps taken to solve a task and is able to solve longer-horizon tasks, validating the hypothesis that explicit state-tracking and goal-tracking help with solving longer horizon reasoning and planning tasks.
+At the core of an LLM agent is the base agent, such as Act (Huang et al., 2022a), ReAct (Yao et al., 2023c), and AdaPlanner (Sun et al., 2023).
+Existing efforts to improve LLM agents build on top of base agents and are usually quite resource-intensive: Wu et al.
+(2024) require human expert annotations of rules; Sun et al.
+(2023) require a code execution environment with carefully crafted code-based prompts; Fu et al.
+(2024) use additional training data together with retrieval augmented generation (RAG) to help the AI agent; Yang et al.
+(2024) use additional training data to fine-tune the LLM; Shinn et al.
+(2023) and Prasad et al.
+(2024) use test-time computation to produce better results.
+Notably, most of the current state-of-the-art methods use ReAct as the base agent.
+In our work, we propose a new base agent, called StateAct.
+Starting with the observation that: i) LLM agents fail to follow the original instruction and goal in longer interactions; and ii) LLMs struggle with long-context despite longer available contexts (Li et al., 2023b; Coelho et al., 2024), we propose two main contributions for improving the base agent.
+To address i), we propose a mechanism for the agent to ‘self-prompt’ at every turn of the interaction to improve staying on track with the main goal.
+Our agent ‘reminds’ itself of the goal at every turn.
+To address ii), we propose ‘chain-of-states’, an extension of chain-of-thought based on state tracking to help the agent stay on track with the current interaction and context.
+The agent keeps track of its state in the environment (such as location and inventory).
+Our experiments show that StateAct achieves near state-of-the-art performance without using additional training data or additional tools, and significantly outperforms ReAct across multiple tasks and eight frontier LLMs of varying sizes.
+Specifically, StateAct improves performance over ReAct by more than 10% on Alfworld (Shridhar et al., 2021), 30% on Textcraft (Prasad et al., 2024), and 7% on Webshop (Yao et al., 2023a).
+Additionally, we validate that StateAct can serve as a drop-in replacement for existing extension methods.
+Using test-time computation (Prasad et al., 2024), we achieve a further 12% performance gain with StateAct on Textcraft using ADaPT.
 2 Background
-AI agents have historically used reinforcement learning (RL) to solve tasks Sutton and Barto (2018).
-With the capabilities of LLMs improving, recent work has combined LLMs and RL, training additional policies or value functions to make predictions (Li et al., 2022; Nottingham et al., 2023).
-2.1 In-context learning approaches
+Figure 1: Overview of LLM-agent methods and their components.
+Due to recent advances, LLMs are now being used as autonomous agents in interactive environments as an alternative to traditional reinforcement learning (RL) Sutton and Barto (2018); Yao et al.
+(2023c); Li et al.
+(2022); Nottingham et al.
+(2023).
+LLM agents now tackle tasks in simulated environments such as Alfworld, Webshop, and Textcraft.
+LLM agents consist of the base agent and a possible extensions, such as fine-tuning (FT), retrieval augmented generation (RAG), test-time scaling (TTS), repeated attempts (REPS) or tools; see Table 1.
+StateAct is a base agent that can be used in combination with extensions.
+2.1 Base agents
 Huang et al.
-(2022a, b) and Singh et al.
-(2022) were among the first to use LLMs directly to act in an interactive environment; their method produces agent actions as output after receiving environment observations as input.
+(2022a, b) were among the first to use LLMs directly to act in an interactive environment; their method produces agent actions as output after receiving environment observations as input.
 ReAct Yao et al.
-(2023b) took this work further by combining ‘acting‘ Huang et al.
-(2022a) and ‘chain-of-thought‘ Wei et al.
+(2023c) take this work further by combining acting Huang et al.
+(2022a) and chain-of-thought Wei et al.
 (2023).
-ReAct establishes state-of-the-art for for approaches based only on in-context learning and while it is a very scalable method, the performance (i.e.
-success rate) is still limited.
+Progprompt and AdaPlanner (Singh et al., 2022; Sun et al., 2023) use code-based prompts to interact with environments.
+Notably, ReAct is the base agent of modern state-of-the-art approaches.
+Our method, StateAct, falls in the category of base agents and is therefore most comparable to ReAct.
+2.2 Extensions of base agents: RAG, fine-tuning, test-time-scaling
 ExpeL Zhao et al.
-(2023) uses additional training data to generate ‘insights’ and ‘success trajectories’ during training.
-At inference time, they look up the closest ‘success trajectories’ as few-shot examples to the agent (as opposed to the fixed few-shot examples that we use) and augment them with these ‘insights’ to perform the final inference.
-They achieve 59% on Alfworld using retrieved ‘success trajectories’ (as few-shot examples) + ‘insights’ and 50% using the same few-shot examples as ReAct + ‘insights’.
-2.1.1 In-context learning and additional tools
-The current state-of-the-art for in-context learning-based approaches in combination with additional tools is ProgPrompt (Singh et al., 2022) and follow-on work AdaPlanner Sun et al.
-(2023).
-AdaPlanner introduces a code-based prompt Li et al.
+(2023) extend ReAct by using additional training data to generate success trajectories during training.
+At inference time, they look up the closest success trajectories as few-shot examples to the agent.
+Follow-on work, AutoGuide Fu et al.
+(2024), uses ReAct as the base agent with additional training data to create state-aware text-based guidelines.
+This ‘knowledge’ of the environment is then used with retrieval augmented generation (RAG) to guide the decision-making process.
+While AutoGuide achieves the best result among RAG approaches, the complexity of the setup makes it less scalable in practice.
+Chen et al.
+(2023); Yao et al.
+(2023c) introduce fine-tuning of ReAct with marginal improvements.
+KnowAgent (Zhu et al., 2025) compile knowledge of the environment and distill this into the LLM to produce better results.
+The best approach, ActRe Yang et al.
+(2024), achieves successful fine-tuning of ReAct by annotating successful trajectories with CoT before fine-tuning.
+Fine-tuning yields the best performance for a specific task, however, requires additional training, which is costly, and does not allow for generalisation.
+ADaPT (Prasad et al., 2024) and THREAD (Schroeder et al., 2024) use test-time computation to achieve better results.
+This approach is in line with works such as tree-of-thought (Yao et al., 2023b), where the AI agent proposes several thoughts or actions in one go during test-time inference and another model (often the same LLM) evaluates these; the top k promising thoughts or actions are then expanded upon in a beam-search manner.
+These methods produce strong results and are very easy to setup, albeit they often require more compute budget compared to base agents.
+In our work we combine ADaPT with StateAct.
+2.3 Alternative methods: tools, hand-crafted rules, multi-agent
+ProgPrompt (Singh et al., 2022) and follow-on work AdaPlanner (Sun et al., 2023) introduce a code-based prompts Li et al.
 (2023a).
-They use code-execution as an additional tool, by executing LLM generated code and feeding the results into the next LLM generation.
-The shortcoming of AdaPlanner is that it requires very complex human-crafted prompts that are hard to scale to new environments and it requires the additional step of code-execution.
-2.2 State tracking in LLM-based agents
-Concurrent work to ours, AutoGuide Fu et al.
-(2024), uses ReAct as the base agent as well as additional training data to create ‘state-aware’ text-based guidelines for the LLM-agent; they then use a type of retrieval augmented generation (RAG) process to guide the decision making process.
-They embed the current observation as a summary (which they call ‘state-aware’) and use an LLM to ‘look up’ the relevant ‘state-aware’ guideline, which is then fed into a final LLM to generate an action.
-Using training data and LLM-based RAG, they achieve 79.1% on top of a ReAct agent on Alfworld.
-Their training and RAG approach could be used complementary to our StateAct LLM-agent [^2] However, the cost of using RAG at every generation step could be significantly higher than our method..
-Chen et al.
-(2024) propose state-tracking as a way to help the agent solve the task without training data.
-Their method differs to ours two-fold.
-Firstly, they employ a complex sequence of components working together (an LLM-based attention over the observations, an LLM-based compilation of a complex state and, finally, a prediction of a program).
-Secondly, their system involves execution of actual programs.
-Our method, on the other hand, requires a straightforward extension of ‘chain-of-thought’ and uses a single LLM call to produce the state, thought and action.
-Additionally, we do not require program execution.
-2.2.1 Fine-tuning approaches
-Previous fine-tuning approaches did not significantly enhance performance Zhou et al.
-(2024); Yao et al.
-(2023b); Shridhar et al.
-(2021).
-A concurrent work, however, ActRe Yang et al.
-(2024b) achieves 83% by fine-tuning gpt-3.5-1106 on additional training data.
-2.3 “Multi-Agent Conversation” approaches
-A new trend is to use multiple LLMs concurrently to ‘chat’ to one another to produce a result. A recent work in this direction by Wu et al. (2023) achieves 67% on Alfworld.
-2.4 Joined rule and LLM-based agents
+They use code-execution as an additional tool, by executing LLM-generated code and feeding the results into the next LLM generation.
+The shortcoming of such code-based prompts is that they require human experts to annotate very long prompts.
+This can be hard to scale to new environments and requires an additional step of code-execution.
 StateFlow Wu et al.
@@
 These FSMs are human expert-crafted states, transitions and rule-based heuristics, where the LLM is asked to perform limited tasks in each of the given states.
-While their performance is 82% on Alfworld, we believe this result is more comparable to a rule-based ‘expert’ than an LLM-based agent (notably close to 100% can be achieved on Alfworld using hand-crafted ‘experts’ Shridhar et al.
-(2021)).
-2.5 Approaches with multiple trials
-Works such as Reflexion Shinn et al.
-(2023) or work by Sun et al.
-(2023); Yang et al.
-(2024a) are methods that perform several evaluations on the same test environments and perform learning loops (during evaluation) on these test-set environments to improve the score.
-While the score necessarily goes up (as only the failed environments are re-run), we do not dive into this as it does not compare to the actual agent performance; additionally, ‘reflection‘ can be used on top of most LLM-based agents (e.g.
-Reflexion uses ReAct as the base agent).
+While this method can achieve a very high score, it is limited, as it requires a human expert to design the FSM.
+Another approach is to use multiple LLMs to ‘chat’ to one another to produce a result. Multi-agent frameworks Wu et al. (2023) only achieve minor improvements over using a single agent.
+2.4 State tracking in LLM-based agents
+Chen et al.
+(2024) propose state-tracking as a way to help the agent solve the task without training data.
+Their method differs from StateAct two-fold.
+Firstly, they employ a complex sequence of components working together, which are an LLM-based attention over the observations, an LLM-based compilation of a complex state and a prediction of a program.
+Secondly, their system involves execution of actual programs.
+StateAct, on the other hand, requires a straightforward extension of CoT and uses a single LLM call to produce the state, thought and action.
+Additionally, it does not require program execution.
+STATLER (Yoneda et al., 2024) also introduce state-tracking for LLM agents.
+While the state has similarity to StateAct, there are notable differences.
+Firstly, Statler is aimed at lower level robotics execution and works with domain specific functions.
+Secondly, Statler produces and requires code to update and read from the state.
+This complex construction of the state is hard to scale to new environments.
 3 Method
-Figure 1: StateAct interacts with the environment directly, without additional tools or resources (a). A skeleton of the StateAct prompt (b).
-StateAct is an LLM-based AI agent that works on top of pre-trained large language models.
-StateAct takes the textual ‘observation’ from the environment and, after a single call to the pre-trained LLM, returns the ‘action’ back to the environment (without the use of additional tools or resources), see Figure 1(a).
+StateAct is an LLM-based AI agent that works on top of pre-trained LLMs.
+It takes the textual ‘observation’ from the environment and, after a single call to the pre-trained LLM, returns the ‘action’ back to the environment, without the use of additional tools or resources.
 StateAct utilises in-context learning Brown et al.
@@
 (2023) to make the agent interact with the environment.
-At the core of the approach is a prompt that consists of few-shot examples of successful ‘interaction traces’ as well as the current interaction trace up to the current step in the environment, see Figure 1(b).
+At the core of the approach is a prompt that consists of few-shot examples of successful interaction traces as well as the current interaction trace up to the current step in the environment.
 An interaction trace consists of alternating observations from the environment and desired (or actual) outputs from the LLM.
@@
 The action is then extracted and passed to the environment to produce the next observation, see Figure 2.
-For example, a sample of the trace for step n in Alfworld is:
+This renders StateAct similar to ReAct and therefore an easy replacement for extension methods.
+3.1 Self prompting
+One of the key ideas of StateAct to overcome the ‘haystack’ challenge for long prompts and therefore long horizon problems (Coelho et al., 2024) is to introduce a mechanism for the LLM to pass an instruction to itself.
+This approach can be applied in many settings and to alleviate various problems (including goal reminding and formatting).
+In our setting, we focus on keeping the Agent on track with the main goal and so the LLM reminds itself of this goal at every turn.
+3.2 Chain of states as state-tracking
+The idea of state-tracking is to introduce ‘structured thoughts’ into the reasoning part of the LLM Agent, specifically by giving the agent small intermediate predictions that can be inferred from the environment and actions.
+This method is different to existing methods such as ReAct (Yao et al., 2023c), where CoT is taken to mean verbal ‘thoughts’.
+The inspiration comes from the original CoT paper (Wei et al., 2023) where the LLM is tasked with producing intermediate calculations.
+For StateAct, the LLM agent is tasked with predicting very specific intermediate steps, such as the current location or the inventory of the agent.
+See Appendix K for a formalization of StateAct.
+Figure 2: Example of StateAct outputs in Alfworld.
+3.3 An example trace
+An example of the trace for step n in Alfworld is:
 Observation n:
-On the countertop 1, you see a tomato 1.
+On the countertop 2, you see a tomato 1.
 LLM Output n:
 >Goal: Put a clean tomato in fridge
-Current Location: countertop 1
+Current Location: countertop 2
 Current Inventory: None
-Thought: I find the tomato, now I need to take it.
+Locations Visited: table 1, countertop 1 Thought: I find the tomato, now I need to take it.
 Action: take tomato 1
-In the above example, the agent produces the action ‘take tomato 1’ in response to the observation ‘On the countertop 1, you see a tomato 1’.
-Figure 2: An example of how StateAct is working. The depicted traces are from the Alfworld environment.
-3.1 Goal-reminders and state-tracking using LLMs
-In order to make StateAct more precise, we introduce goal and state tracking.
-We denote $\pi$ as the policy of an AI-agent, in the standard case at time step $t$, the policy predicts action $a_{t}$, given the history of observations and actions $[o_{t},a_{t-1},...,a_{0},o_{0}]$.
-$\pi(a_{t}|o_{t},a_{t-1},...,a_{0},o_{0})$ (1)
-Where $a_{t}$ is the action produced by the agent at time step t and $o_{t}$ is the observation produced by the environment at time step t after receiving action $a_{t}$ as input.
-Usually, the first observation $o_{0}$ also contains the ‘goal’ description for the given environment.
-For our case, we need to enhance the policy to incorporate the ‘state’.
-Similar to previous work Yao et al.
-(2023b), we introduce the context vector, $c_{t}$.
-The context vector contains the action, as well as the additional predictions of the agent , i.e.
-$c_{t}=(g_{0},s_{t},r_{t},a_{t})$.
-Where $g_{0}$ is the goal and it always remains the same (for a given environment) and uses the goal extracted from $o_{0}$; $s_{t}$ represents the predicted state at time step $t$; $r_{t}$ represents ‘chain-of-thought’-style ‘reasoning’ at time step $t$; and $a_{t}$ represents the action at time step $t$, as before.
-The new policy $\pi$ then becomes:
-$\pi_{contextual}(c_{t}|o_{t},c_{t-1},...,c_{0},o_{0})$ (2)
-In our case, the LLM acts as $\pi_{contextual}$ and produces the context vector at every time step.
+In this example, the agent produces the action ‘take tomato 1’ in response to the observation ‘On the countertop 2, you see a tomato 1’.
 4 Experimental setup
-Our aim is to study the long-range acting, planning and reasoning capabilities of LLM-based agents.
-To achieve this, in line with previous work, we turn to simulated environments as an evaluation framework and to API-based state-of-the-art large language models.
-Specifically, we use Alfworld Shridhar et al.
-(2021), a household robotics environment, and Webshop Yao et al.
-(2023a), an online shopping environment, as our simulated environments.
-In line with previous work, we use OpenAI’s [^3] https://openai.com, last accessed June 2024.
-gpt-3.5-turbo-1106.
-We, also, evaluate Alfworld on additional LLMs and validate our results.
+Our aim is to study reasoning capabilities of LLM-based agents.
+In line with previous work, we use Alfworld Shridhar et al.
+(2021), a household robotics environment; Webshop Yao et al.
+(2023a), an online shopping environment; and Texcraft (Prasad et al., 2024), a text-based tool crafting game, as our evaluation environments.
+We mainly evaluate StateAct as the base agent.
+However, we also show that StateAct can be used as a drop-in replacement for ReAct and validate it on test-time scaling.
+RAG (AutoGuide) and fine-tuning (ActRe)-based extensions are left for future work, as these are expensive and complex to run due to training on additional data and use of additional tools.
 4.1 Alfworld
@@
 Alfworld Shridhar et al.
-(2021) is based on a 3D, visual, household robotics environment called Alfred Shridhar et al.
+(2021) is based on a 3D visual household robotics environment called Alfred Shridhar et al.
 (2020), which was translated into a text-based environment for ease of use for language-based AI models, see Figure 3.
@@
 clean an apple and put it on a table.
-The robot (or agent) then needs to perform a series of ‘high-level’ operations to accomplish the tasks, e.g.
+The robot (or agent) then needs to perform a series of high-level operations to accomplish the tasks, e.g.
 ‘go to fridge 1’, ‘open fridge 1’.
-At every step, the environment provides either a textual observation or feedback that the command has failed, e.g.
+At every step, the environment provides either a textual observation or the feedback that the command has failed, e.g.
 ‘You open the fridge 1’, ‘You see apple 1’.
 The underlying text engine is based on Textworld Côté et al.
-(2019), see Appendix A for a complete list of the commands and further details on the environments.
-4.1.1 Alfworld correction
-In our research, we identified that Alfworld has a specific syntactic feature for the put command, namely put <object> in/on <place>, where “in/on” needs to be written exactly this way.
-Using only “in” or only “on” produces a failed command.
-We observed this issue with LLMs in this environment and we propose a simple fix for it.
-We map: 1.
-“put <object> in <place>” and 2.
-“put <object> on <place>” to the command accepted by Alfworld, namely “put <object> in/on <place>”.
-Methods such as AdaPlanner Sun et al.
-(2023) have avoided this issue because they use code-based prompts and regex parsers.
-However, methods such as ReAct Yao et al.
-(2023b) and ExpeL Zhao et al.
-(2023) have been affected by it therefore lowering their potential performance.
-We report results for ReAct and StateAct both with and without these corrections.
+(2019), see Appendix A for a complete list of the commands and details of the environments.
 4.2 Webshop
@@
 In line with previous work, we use the text-based version of Webshop, where all descriptions of the website are given in text form, see Figure 4.
-Webshop features a realistic large-scale product catalogue, a search engine and very varied product attributes depending on the category of product, see Appendix B for more details.
-In total, the test set consists of 500 examples and each one is of the type “search and buy a product”.
-Overall, Webshop has a maximum of 15 steps and two commands: 1.
+Webshop features a realistic large-scale product catalogue, a search engine, and very varied product attributes depending on the category of product, see Appendix B for more details.
+In total, the test set consists of 100 examples and each one is of the type “search and buy a product”.
+Overall, Webshop has a maximum of fifteen steps and two commands: 1.
 search[<query>], 2.
 click[<button>].
-4.3 In context learning
+4.3 Textcraft
+Textcraft Prasad et al.
+(2024) is an environment based on the popular game Minecraft, where the task of the agent is to craft items.
+The environment is fully text-based.
+We use prompts and implementations based on the ADaPT paper (Prasad et al., 2024) that introduced this environment.
+We also use this environment to analyse whether StateAct performs well in combination with other methods such as ADaPT.
+See Appendix C for more details.
+4.4 In-context learning
 Since ReAct Yao et al.
-(2023b) forms the underlying agent for many current Zhao et al.
+(2023c) forms the underlying agent for many current Zhao et al.
 (2023) and state-of-the-art approaches Fu et al.
-(2024), we use the same few-shot ‘interaction traces’ as ReAct.
-The main reason is to have a fair comparison and to isolate additional effects - such as performance change - from different in-context examples.
-In total, Alfworld has six types of tasks and ReAct uses two in-context examples per task type to prompt the language models.
+(2024); Yang et al.
+(2024); Prasad et al.
+(2024), we use the same few-shot interaction traces as ReAct.
+The main reason is to have a fair comparison and to isolate additional effects, such as performance change, from different in-context examples.
+Alfworld, for example, has six types of tasks and ReAct uses two in-context examples per task type to prompt the language models.
 On average, each ReAct example ranges from 352 words to 591 words (590 tokens to 935 tokens).
-For our study, we reuse the observations, thoughts and actions, and annotate these examples further with goal ‘reminders’ and state tracking, which results in a range from 484 to 911 words (807 tokens to 1458 tokens) per example.
+For our study, we reuse the observations, thoughts and actions, and annotate these examples further with the goal and state, which results in a range from 484 to 911 words (807 tokens to 1458 tokens) per example.
 During our annotation, we discovered minor errors in the ReAct prompts and fixed them as well.
-We release all our annotations with our code release.
 In comparison, AdaPlanner Sun et al.
 (2023), uses a different code-based approach and the prompt has 1104 words (2015 tokens) on average.
-We use the two-shot examples from ReAct for Alfworld and the one-shot example from ReAct for Webshop in all our experiments.
+We use the two-shot examples from ReAct for Alfworld, the one-shot example from ReAct for Webshop and the few-shot prompts of ReAct from ADaPT for Textcraft in all our experiments.
 4.5 Models
-In line with previous work, we focus our main experiments are based on the API-based LLMs to compare performance fairly.
-Concretely, ‘gpt-3.5’ level models were used in previous work including current state-of-the-art ActRe (Yang et al., 2024b) and state-of-the-art for in-context learning ReAct (Yao et al., 2023b).
-Many OpenAI models have become deprecated; notably, all models from ReAct and AdaPlanner Sun et al.
-(2023) davinci-002, gpt-3.5-turbo-0301, and gpt-3.5-turbo-0613 are now deprecated.
-Therefore, we re-implemented ReAct and ran the experiments to determine the most suitable model, see Table 1.
-We establish that gpt-3.5-turbo-1106 is the best performing (from the ones that remain available) on ReAct and we therefore chose this model for our main experiments.
-In our work, we also compare our method using newer state-of-the-art models to show that our method generalises to new models.
-Specificially, we use the gpt-4o-mini [^4] Again, we chose the smaller model due to cost.
-and Mixtral (Jiang et al., 2024) models.
-We note that full GPT-4 level models are prohibitively expensive [^5] A single evaluation run on Alfworld costs approx.
-\$8 using gpt-3.5 and ReAct, whilst gpt-4 costs 10+ times more..
-We use temperature 0 for all experiments and sample only the top 1 response, see Appendix C for the exact settings.
-Method Version AW WS
-ReAct gpt-3.5-turbo-1106 40.7 18.2
-ReAct gpt-3.5-turbo-0125 30.37 14.6
-Table 1: ReAct success rate on Alfworld (AW) and Webshop (WS), using available gpt-3.5 models.
-4.5 Metrics
-In terms of metrics, we use the pre-defined metrics of Alfworld and Webshop, namely success rate (SR).
-Success is a binary metric per each environment in the respective test sets (134 for Alfworld and 500 for Webshop respectively).
-Success in Alfworld means the agent has successfully completed the whole task.
-In Webshop, it means the agent has bought an item that has a one hundred percent match with the desired item based on a partially hidden list of shopping item attributes (e.g.
-the colour, size, price, etc.).
+In our work, we compare our method using newer state-of-the-art models, architectures and sizes to show that our method generalises.
+Specifically, we use Mistral-Small-24B-Instruct (Jiang et al., 2023), Qwen-2.5-7B,14B,32B-Instruct (Qwen et al., 2025) and Gemma2-27B-Instruct (Team et al., 2024).
+We note that api-based models such as OpenAI’s models are generally very expensive [^2] A single evaluation run on Alfworld costs approx.
+\$8 using gpt-3.5 and ReAct, whilst gpt-4 costs 10+ times more.
+and lack rigorous reproducibility standards [^3] Since we do not have access to weights and inference settings and models become regularly deprecated..
+Nonetheless, we include experiments using gpt-3.5 and gpt-4o-mini.
+We use temperature 0 for all experiments and sample only the top 1 response; we use vllm for inference (Kwon et al., 2023), see Appendix D & E for details.
 5 Results
-5.1 Alfworld
-For Alfworld, we present the results for ReAct, AdaPlanner with and without code execution, and StateAct (ours), which consists of goal + state + thought + action.
-We also show StateAct without each of the components (i.e.
-without goal, state and thought).
-Interestingly, we find that, contrary to previous findings, ‘thought’ or ‘reasoning’ actually sometimes hinders the performance.
-In Table 2, we can see that StateAct with all goal + state + thought and the correction performs the best.
-It outperforms ReAct with correction by around 13 points (when using the same GPT model for ReAct) and by around 9 points (when using the better model for ReAct).
-StateAct also outperforms ReAct by 22 points when corrections are not used.
-Furthermore, StateAct even outperforms AdaPlanner by 2.48 points, an approach that uses regex for command-mapping (similar to our correction) along with using code-execution.
-Perhaps the most surprising finding is that the simple correction described in Section 4.1.1 leads to a 16- and 23-point jump for ReAct and a 27-point jump for StateAct.
-This indicates that the model generally performs very well; however, it struggles with minute differences in domain-specific syntax.
-Method GPT-3.5 SR %
-Baselines w/o corrections
-ReAct 0301* 51.9
-ReAct 1106 40.7
-ReAct 0125 30.37
-ReAct (joined***) 1106 36.30
-Results w/o corrections
-StateAct (ours) 1106 50.37
-- w/o thought 1106 62.96
-- w/o goal 1106 42.96
-- w/o state 1106 44.44
-Baselines with corrections
-ReAct + corr. 0301* 68.15
-ReAct + corr. 1106 63.70
-AdaPlanner No-Exec** 0301* 46.66
-AdaPlanner 0301* 75.56
-Results with corrections
-StateAct + corr. (ours) 1106 77.04
-- w/o thought 1106 64.44
-- w/o goal 1106 74.04
-- w/o state 1106 64.44
-Table 2: Success Rate (SR) on the 134 test-set examples from Alfworld.
-*gpt-3.5-0301 was deprecated in June 2024.
-**No-Exec means AdaPlanner without code execution.
-***joined means that thought + action are produced at every turn.
-AdaPlanner results are from Sun et al.
-(2023).
-All other results are single run.
-5.2 Webshop
-For Webshop, we present the results for ReAct and StateAct.
-Similarly to Alfworld, we also present the results of StateAct without each of goal, state and thought, see Table 3.
-Interestingly, we see that removing thought produces the highest results and outperforms ReAct by 10 points.
-Our hypothesis is that domain-specific syntax, which is more prevalent in Webshop than Alfworld, conflicts with using verbose thoughts.
-Method GPT-3.5 Success Rate %
-Baselines
-Rule-based* N/A 9.60
-RL* N/A 17.60
-ReAct 1106 17.80
-Results
-StateAct (ours) 1106 17.00
-- w/o thought 1106 27.80
-- w/o goal 1106 20.40
-- w/o state 1106 21.00
-Table 3: Success Rate (SR) on the 500 test-set examples from Webshop.
-*results taken from Yao et al.
-(2023a).
-For StateAct without thought and ReAct Results are the average of two runs.
-Other Results are single run.
-5.3 Results using additional models
-We validate our results further with additional LLMs.
-Again we see a significant boost in performance, see Table 4.
-We experimented with gpt-4o-mini and Mixtral [^6] The concrete version of Mixtral was ‘mixtral-8x22b-instruct-v0.1’.
-(Jiang et al., 2024) on Alfworld.
-ReAct achieves 68.15 and 72.59 respectively, while StateAct achieves 71.85(+3.7) and 83.70(+11.2).
-Method Model SR %
-Baselines
-ReAct Gpt-4o-mini 68.15
-ReAct Mixtral-8x22B* 72.59
-Results
-StateAct (ours) Gpt-4o-mini 71.85
-StateAct (ours) Mixtral-8x22B* 83.70
-Table 4: Success Rate (SR) on the 134 test-set examples from Alfworld. Results are ‘single run’. *=The concrete version of Mixtral is ‘mixtral-8x22b-instruct-v0.1’.
+Overall we present results of state-of-the-art methods such as ActRe and ADaPT, as well as the underlying base agent ReAct.
+In Table 1 we see that while methods that rely on ReAct + extensions, such as ActRe and AutoGuide outperform StateAct overall, they also rely on additional training data and computation.
+Furthermore, StateAct achieves the best result among base agents, outperforming ReAct between 7% and 30% and achieves comparable result with the state-of-the-art methods, while not using any additional tools or data.
+For example, AutoGuide (ReAct + RAG) gets 0.79, while StateAct gets 0.77.
+Method Score
+State-of-the-art methods
+AdaPlanner (Code-prompt + exec.)1 0.75
+AutoGuide (ReAct + RAG)2 0.79
+ActRe (ReAct + fine-tuning)3 0.83
+ADaPT (ReAct + test time scaling)4 0.72
+Base agents
+Act (few-shot only) 0.41
+ReAct (few-shot only) 0.64
+AdaPlanner (Code-prompt only)1 0.45
+StateAct (ours, few-shot only) 0.77
+Table 1: Results on the 134 test samples of Alfworld using gpt-3.5.
+ReAct and StateAct scores are single run with greedy decoding and gpt-3.5-1106.
+1=code-execution (Sun et al., 2023), 2=(Fu et al., 2024), 3=(Yang et al., 2024), 4=(Prasad et al., 2024).
+5.1 Base agent comparison
+Agent Name Mistral-24B Qwen-7B Qwen-14B Qwen-32B Gemma-27B Average $\uparrow$
+Alfworld*
+ReAct 0.44 0.10 0.75 0.89 0.71 0.58
+StateAct 0.49 0.46 0.78 0.90 0.68 0.66
+Webshop**
+ReAct 0.34 0.19 0.22 0.27 0.26 0.26
+StateAct 0.35 0.12 0.33 0.32 0.29 0.28
+Textcraft***
+ReAct 0.33 0.02 0.31 0.31 0.18 0.23
+StateAct 0.40 0.04 0.37 0.40 0.34 0.31
+Table 2: Base agent performance across different models and environments.
+*=134 Test Environments from Alfworld.
+**=100 Test Environments from Webshop.
+***=100 Test Environments from Textcraft.
+M=Mistral-Instruct-2501, Q=Qwen2.5-Instruct, G=Gemma 2-Instruct.
+We use greedy decoding (temperature=0).
+Since ReAct is the previous best base agent and forms the basis of state-of-the-art approaches, we compare against it in detail.
+The results in Table 2 show that StateAct outperforms ReAct across three different benchmarks and five different models.
+Sometimes the difference is quite substantial, with StateAct outperforming ReAct by more than 10 points.
+Across all 15 experiments only in two ReAct performs better than StateAct: in Webshop, where the 7B model is likely overwhelmed by the amount of textual input it receives, as Webshop has a lot of verbal input; and in the case of Gemma2-27B ReAct slightly outperforms StateAct on Alfworld; a hypothesis is that Gemma has a limited context length of 8192, while Alfworld requires long traces due to the longer step length.
+We further validate our results with additional LLMs that are too large to fit on a single GPU or are closed source.
+In Table 3 we see a significant performance increase.
+Using gpt-3.5, gpt-4o-mini and Mixtral-8x22B [^4] Mixtral-8x22b-instruct-v0.1 was queried using Nvidia’s NIM API https://developer.nvidia.com/nim [Last Accessed March 2025].
+(Jiang et al., 2024) on Alfworld, ReAct achieves 63.7, 68.15 and 72.59, while StateAct achieves 77.04 (+13.3), 71.85(+3.7) and 83.70(+11.2) respectively.
+Method Model Success Rate %
+ReAct Gpt-3.5 0.64
+StateAct Gpt-3.5 0.77
+ReAct Gpt-4o-mini 0.68
+StateAct Gpt-4o-mini 0.72
+ReAct Mixtral-8x22B 0.73
+StateAct Mixtral-8x22B 0.84
+Table 3: Results on the 134 test examples from Alfworld. Results are single run and greedy. Models used: gpt-3.5-1106, gpt-4o-2024-07-18, mixtral-8x22b-instruct-v0.1.
+5.2 Base agents + test time scaling
+An important contribution is to validate that StateAct can be used as a drop-in replacement with advanced methods.
+To this end we validate StateAct using test-time scaling using the ADaPT method.
+Starting with ADaPT, which is based on ReAct, as the starting point(Prasad et al., 2024), we enhance their method using StateAct.
+In Table 4 we can clearly see that StateAct scales well with test time scaling jumping in performance by 39%, and surpassing ADaPT+ReAct by 12%.
+Model/Agent Name Normal +ADaPT
+Mistral-24B
+ReAct 0.33 0.53
+StateAct 0.40 0.64
+Qwen2.5-7B
+ReAct 0.02 0.09
+StateAct 0.04 0.11
+Qwen2.5-14B
+ReAct 0.31 0.55
+StateAct 0.37 0.53
+Qwen2.5-32B
+ReAct 0.31 0.64
+StateAct 0.40 0.62
+Gemma-2-27B
+ReAct 0.18 0.19
+StateAct 0.34 0.35
+Average
+ReAct 0.23 0.4
+StateAct 0.31 0.45
+Table 4: Comparison of test-time scaling performance on 100 test samples from Textcraft.
+Normal refers to using just the base agent.
++ADaPT means running the respective base agent with test time scaling using the ADaPT method.
+ADaPT code is adapted to run ReAct and StateAct.
+We use greedy decoding and $d_{max}$ = 2.
 5.3 Summary of results
-In conclusion, we found that our simple goal-reminding and state-tracking approach that purely relies on in-context learning outperforms previous in-context learning approaches by more than 10 points and even outperforms leading approaches that rely on code-execution.
-Interestingly, we found that the approaches are quite sensitive to domain-specific syntax and, when this is the case, ‘thoughts’ that are verbose can hinder performance.
+StateAct establishes itself as the best-performing base agent, surpassing ReAct by 7–30% across multiple benchmarks while requiring no additional training data or external tools.
+Although advanced methods like ActRe and AutoGuide achieve higher scores, they rely on costly training and retrieval.
+We also validate StateAct as a drop-in replacement for ReAct in test-time scaling.
+These results highlight the performance and usability of StateAct as a robust foundation for LLM-based agents.
 6 Analysis and Ablations
-In the results section, we discovered that our methods perform better than previous state-of-the-art.
-Our method shows strong performance with in-context learning without resorting to additional tools, data or bigger models.
-In this section, we analyse our results further and also answer our second hypothesis that goal ‘reminding’ and state-tracking help with long-range reasoning.
-For all ablation studies, we focus on Alfworld as it has two favourable properties over Webshop.
+In the results section, we discovered that our methods outperform the previous state-of-the-art base agent, ReAct.
+Our method shows strong performance with in-context learning without resorting to additional tools or data.
+In this section, we analyse our results further and also show that self prompting and state-tracking help with long-range reasoning.
+For most ablation studies, we focus on Alfworld as it has two favourable properties over Webshop and Textcraft.
 Firstly, Alfworld has a longer time horizon (50 steps vs.
-15 in Webshop), with tasks taking, on average, less than 10 steps in Webshop and around 20 to 30 steps in Alfworld.
-Secondly, Alfworld has much less domain-specific syntax and is purely text-based, while Webshop has a more specific syntax to follow.
-6.1 Do goal reminders help with long-range tasks?
-For this purpose, we compare the original ReAct (thought + action) with the goal included, i.e.
+15 in Webshop, 40 in Textcraft).
+Secondly, Alfworld is more realistic than Textcraft, as Alfworld is a robotics environment, while Textcraft is based on a game.
+For completeness, we include ablation results in Table 6 and full results in Appendix L,M,N.
+6.1 Does self-prompting help with long-range tasks?
+For this purpose, we compare the original ReAct (thought + action) with the self-prompting included, i.e.
 StateAct (goal + thought + action).
-In Figure 5 we can see that, while the performance of both ReAct and StateAct goes down as there are more steps, the goal tracking has a substantially better relative performance as the number of steps increases.
-This finding is in-line with our original motivation that LLM agents deteriorate in performance as the prompts and interactions get longer.
+In Figure 5 we can see that, while the performance of both ReAct and StateAct goes down as there are more steps, the goal tracking has better relative performance as the number of steps increases and is able to solve longer tasks of 40 to 50 steps.
+This finding is in line with our original motivation, that LLM agents deteriorate in performance as the prompts and interactions get longer.
 To verify that this actually means that goal tracking helps with performance, as opposed to just increasing the number of steps it takes to solve a task, we calculate the average number of steps for ReAct [^5] We ignore ‘thought’ turns for ReAct as otherwise ReAct would have even more steps.
 and StateAct.
-Table 5 clearly show that ReAct, with an average of 38.84 steps to solve an environment, is the least efficient whilst StateAct, with an average of 28.96 steps to solve an environment, is the most efficient.
-This shows that not only does goal reminders help with longer range tasks, it also helps with efficiency, by shortening the tasks.
-See Appendix E for more discussion.
-Figure 5: Goal vs. No Goal, on the test set of Alfworld, using gpt-3.5-turbo-1106 without correction.
+Table 5 clearly shows that ReAct, with an average of 31.49 steps to solve an environment, is the least efficient whilst StateAct, with an average of 19.11 steps to solve an environment, is the most efficient.
+This shows that not only does self-prompting help with longer range tasks, it also help with efficiency, by shortening the tasks.
+See Appendix G for more discussion.
 6.2 What effects does state-tracking have?
@@
 In fact, we can see that reasoning alone is unable to solve tasks longer than 40 steps, while with both state tracking and goal-tracking longer-range tasks can be solved.
-Also, looking at Table 5 we see that state-tracking makes the model the most efficient [^8] In terms of cost, we found that, despite our method using a twice-longer prompt, our cost remains similar, at around \$8 for the full Alfworld run, since we solve tasks more efficiently and use fewer steps..
-Concretely, ReAct has 24/134 examples that are in the bucket ‘40-50’ and solves 0, while our method has 6/134 examples and solves 4.
-Therefore, we find that explicit state-tracking helps with long-range tasks and enables the agent to solve the tasks more efficiently.
-Figure 6: State vs. No State, on the 134 test examples from Alfworld. gpt-3.5-turbo-1106 with correction.
+Also, looking at Table 5 we see that state-tracking makes the model the most efficient [^6] Despite StateAct using a twice-longer prompt, our cost remains similar to ReAct, at around \$8 for the full Alfworld run, since we solve tasks more efficiently and use fewer steps..
+Concretely, ReAct has 24/134 examples that are in the bucket ‘40-50’ and solves 0, while StateAct has 6/134 examples and solves 4.
+Therefore, we find that explicit state-tracking helps with long-range tasks and to solve the tasks more efficiently.
+Figure 5: State vs. No State, on the 134 test examples from Alfworld, using gpt-3.5-turbo-1106
 Method Avg. Steps $\downarrow$
-ReAct 38.84
-StateAct (goal+thought+action) 31.19
-StateAct (full) 28.96
-ReAct with corr. 31.49
-StateAct with corr. 19.11
+ReAct 31.49
+StateAct (goal, state, thought) 19.11
+- w/o goal (self-prompt) 20.09
+- w/o state 22.50
 - w/o thought 23.76
-- w/o goal 20.09
-Table 5: Average number of Steps (Avg. Steps) [lower is better] on test-set of Alfworld, using gpt-3.5-1106.
+Table 5: Average number of steps (Avg. Steps) [lower is better] on the test set of Alfworld, using gpt-3.5-1106.
 6.3 Does the model perform actual state-tracking?
-We ask ourselves whether the model is actually performing state-tracking.
-For that purpose, we look at Alfworld and construct a self-verification algorithm that is able to track the state heuristically [^9] See Appendix D for details.
+We investigate whether the model is actually performing state-tracking.
+For that purpose, we look at Alfworld and construct a self-verification algorithm that is able to track the state heuristically [^7] See Appendix F for details.
 based on the actions the agent takes.
@@
 We compare the ‘gold’ state against the predicted state.
-Figure 7 shows that StateAct does in fact correct state-tracking 88% of the time.
+Figure 6 shows that StateAct achieves a state-tracking accuracy of 88%.
 We also observe that thoughts and goals help the state-tracking.
 Figure 6: State-tracking accuracy for StateAct on 134 test examples of Alfworld, using gpt-3.5-1106.
-6.4 Does json structure help with performance?
-Since we found that domain-specific syntax hinders performance, we investigated whether adding a structured format like json would help.
-For this purpose, we re-ran StateAct on Alfworld, but translated the state into a json format, see Appendix G for more details.
-Surprisingly, we found that the json format hinders performance significantly, see Table 6.
-However, we also see that corrections help the json format less, indicating that json helps with syntax but hinders performance.
-Method SRSR (+json)%
-StateAct 50.37 45.19 (-5.2)
-StateAct w/ corr. 77.04 58.52(-18.5)
-Table 6: No-json vs. json. Success Rate (SR) on the test set of Alfworld, using gpt-3.5-1106.
-7 Conclusion and future work
-Inspired by the challenge that LLMs and LLM agents struggle with long context, we propose a novel in context learning method for LLM agents - StateAct.
-We introduce ‘chain-of-states’, which relies on in-context learning alone and establish a new state-of-the-art for agents that do not perform training, even against methods that use code-execution.
-StateAct outperforms the previous state-of-the-art that uses in-context learning alone, by between 9% and 20% given different models and tasks, and outperforms in-context learning with tools (i.e.
-code-execution) by 3%, while performing on par with methods that require additional training data.
-In conclusion, we show that explicit state-tracking and goal reminders help the model solve long range reasoning tasks and interestingly also make the model more efficient.
+6.4 Do ‘thoughts’ help?
+While chain-of-thought helps in most cases, surprisingly, we see that in some cases chain-of-thought can be harmful.
+Specifically, in Webshop ‘thoughts’ actually harm overall performance across different agent and model types, see Table 6.
+A hypothesis is the verbosity of the Webshop environment, thus confusing the model with ‘verbal’ thoughts.
+Agent Name AW WS TC
+Baselines
+Act 0.51 0.19 0.27
+Thought+Act (ReAct) 0.58 0.26 0.23
+Our Methods
+State+Act 0.51 0.31 0.22
+State+Thought + Act 0.58 0.18 0.34
+Goal+Act 0.63 0.24 0.26
+Goal+Thought + Act 0.65 0.18 0.30
+Goal+State+Act 0.56 0.29 0.21
+Goal+State+Thought+Act 0.66 0.28 0.31
+Table 6: Ablation table. Averaged results across Mistral, Qwen, Gemma on AW=Alfworld, WS=Webshop and TC=Textcraft. Goal=Self-prompt.
+7 Conclusion
+Our work is driven by the fundamental challenge that LLM agents struggle with long context and keeping on track with instructions.
+The current state-of-the-art to overcome such challenges propose extensions on top of the base agent, ReAct.
+In contrast, we introduced a novel base agent, StateAct, an in-context learning method that leverages chain-of-states and self-prompting to significantly enhance the capabilities of LLM agents.
+By rethinking how agents track and utilize state information, StateAct establishes a new state-of-the-art for base agents, surpassing ReAct by 9% to 20% across different models and tasks.
+Furthermore, we demonstrate that StateAct scales effectively when combined with advanced techniques such as ADaPT, reinforcing its usability.
+Beyond raw performance, our analysis uncovers a crucial insight: StateAct not only improves reasoning but also enhances efficiency, allowing agents to achieve better results with fewer steps.
+This suggests that integrating structured state-tracking and self-prompting cues helps mitigate the well-documented long-context issue in LLMs.
+Our findings suggest that methods like StateAct can serve as a practical and efficient way to improve LLM-based agents without requiring costly external augmentations or data such as retrieval or fine-tuning.
+By enabling agents to manage their own state explicitly, StateAct provides a scalable approach to improving reasoning and decision-making across a range of tasks.
+This makes StateAct a viable drop-in solution for current systems, offering both better performance and greater efficiency in LLM agent tasks.
 8 Ethical Considerations
@@
 We should consider the environment and financial resources for reproduciblity of our work.
-We aimed to address this concern by using gpt-3.5-turbo level models, reporting costs and minimising the cost of our method.
+We aimed to address this concern by models that are less computationally demanding such as gpt-3.5-turbo level models or open source models that fit on a single GPU (A100, 80GB), reporting costs and minimising the cost of our method.
 8.2 Hallucinations in LLMs
@@
 9.1 Languages and evaluation benchmarks
-We evaluated our method only in the English language and on two evaluation benchmarks.
-While we do not expect major changes in other languages, this is something that should be investigated.
-Furthermore, performance on other benchmarks should be evaluated as well.
+We evaluated our method only in the English language and on three evaluation benchmarks. While we do not expect major changes in other languages, this is something that should be investigated.
 9.2 Reasoning traces that rely on human judgement
@@
 Example: use desklamp 1
+A.2.1 Alfworld correction
+In our research, we identified that Alfworld has a specific syntactic feature for the put command, namely put <object> in/on <place>, where “in/on” needs to be written exactly this way.
+Using only “in” or only “on” produces a failed command.
+We observed this issue with LLMs in this environment and we propose a simple fix for it.
+We map: 1.
+“put <object> in <place>” and 2.
+“put <object> on <place>” to the command accepted by Alfworld, namely “put <object> in/on <place>”.
+Note: In the latest release of Alfworld (December 2024) this was fixed by replaceing the put <obj> in/on <place> command with move <obj> to <place> command.
+In our work we report results on the latest version of alfworld.
 A.3 License
@@
 Webshop has the permissible Princeton license; we used it in line with the license.
-Appendix C Code snippet to call OpenAI / GPT-3.5
+Appendix C Textcraft
+C.1 Commands and environment
+Textcraft has one environment type: ‘craft’, as well as three commands: 1. inventory, 2. craft, 3. get
+C.2 Crafting Recipes
+Textcraft has crafting recipes that range from easy to hard. Where hardness is measured by the ‘depth’ of the crating recipe. Specifically, depths of 2, 3 and 4 are present in the dataset.
+C.3 License
+Textcraft is published under the permissible MIT license.
+Appendix D Compute Requirements for local LLMs
+The exact code will be released upon publication.
+However, to help reproducibility we ran all experiments on single A100 80GB GPUs.
+In terms of software we used: vLLM for inference.
+The hyper-parameters were set to: max model length 16000 (except for Gemini, where we used 8192), temperature = 0, datatype="auto" (which results in bfloat16).
+Appendix E Code
+E.1 Code Snippet to call Local LLMs
+from vllm import LLM, SamplingParams
+self.llm = LLM(
+model=model,
+tensor_parallel_size=tensor_parallel_size,
+gpu_memory_utilization=0.95,
+max_model_len=max_model_len,
+dtype="auto"
+messages = [
+# {"role": "system", "content": "You are a helpful assistant."},
+{"role": "user", "content": prompt[-self.max_model_len:]}
+text = self.tokenizer.apply_chat_template(
+messages,
+tokenize=False,
+add_generation_prompt=True
+sampling_params = SamplingParams(
+temperature=self.temperature,
+top_p=1.0,
+repetition_penalty=1.00,
+max_tokens=min(2000,self.max_model_len),
+stop = self.stop_sequences,
+seed = self.seed
+outputs = self.llm.generate([text], sampling_params)
+return outputs[0].outputs[0].text
+E.2 Code Snippet to call OpenAI / GPT-3.5
 client = openai.OpenAI(
@@
 […]
-A prompt is given in Appendix H.
-Appendix D Heuristic State-tracking Explained
+A prompt is given in Appendix J.
+Appendix F Heuristic State-tracking Explained
 Heuristic state tracking is based on the idea that the state can be inferred automatically if one follows the actions of the agent and observations of the environments.
@@
 This rule based ‘state-tracking’ is how the heuristics work.
-Appendix E ‘Step Length Analysis’ Discussion
+Appendix G ‘Step Length Analysis’ Discussion
 An alternative to calculating and comparing step length could be ‘gold solutions’ to measure optimal step length and optimality of an agent.
@@
 Thus a ‘difficulty’ measure would be needed instead, but it is ambiguous.
-Appendix F Potential Future Work Directions
+Appendix H Potential Future Work Directions
 We found that ‘thoughts’ or explicit reasoning do not always help performance.
@@
 Finally, problems related to domain-specific syntax are also an interesting avenue for future work.
-Appendix G StateAct Json Format
+Appendix I Does JSON structure help StateAct performance?
+We also investigated whether adding a structured format like json would help.
+For this purpose, we re-ran StateAct on Alfworld, but translated the state into a json format, see section I.1 for more details.
+Surprisingly, we found that the json format hinders performance, see Table 7.
+Method SRSR (+json)%
+ReAct 63.70 62.96(-0.74)
+StateAct (complete) 77.04 58.52(-18.5)
+Table 7: No-json vs. json. Success Rate (SR) on the test set of Alfworld, using gpt-3.5-1106.
+I.1 JSON prompt
 We translate the text-based StateAct prompt:
@@
 "action": "go to fridge 1"
-Appendix H StateAct Prompts
-H.1 Example Alfworld prompt for the ‘heat’ task
+Appendix J StateAct Prompts
+J.1 Example Alfworld prompt for the ‘heat’ task
 Interact with a household to solve a task.
@@
 <CURRENT TASK>
-H.2 Example Webshop prompt
+J.2 Example Webshop prompt
 Webshop
@@
 Action: click[Buy Now]
+Appendix K Formalizing StateAct
+Let us denote by $\pi$ the policy of an AI-agent, in the standard case at time step $t$, the policy predicts action $a_{t}$, given the history of observations and actions $[o_{t},a_{t-1},...,a_{0},o_{0}]$.
+$\pi(a_{t}|o_{t},a_{t-1},...,a_{0},o_{0})$ (1)
+Where $a_{t}$ is the action produced by the agent at time step t and $o_{t}$ is the observation produced by the environment at time step t after receiving action $a_{t}$ as input.
+Usually, the first observation $o_{0}$ also contains the ‘goal’ description for the given environment.
+For our case, we need to enhance the policy to incorporate the ‘state’.
+Similar to previous work Yao et al.
+(2023c), we introduce the context vector, $c_{t}$.
+The context vector contains the action, as well as the additional predictions of the agent , i.e.
+$c_{t}=(g_{0},s_{t},r_{t},a_{t})$.
+Where $g_{0}$ is the goal and it always remains the same (for a given environment) and uses the goal extracted from $o_{0}$; $s_{t}$ represents the predicted state at time step $t$; $r_{t}$ represents ‘chain-of-thought’-style ‘reasoning’ at time step $t$; and $a_{t}$ represents the action at time step $t$, as before.
+The new policy $\pi$ then becomes:
+$\pi_{contextual}(c_{t}|o_{t},c_{t-1},...,c_{0},o_{0})$ (2)
+In our case, the LLM acts as $\pi_{contextual}$ and produces the context vector at every time step.
+Appendix L Additional results for Alfworld
+For Alfworld we can see that the full StateAct performs best.
+Interstingly, adding state in does not always help model performance.
+A potential explanation for this can be that state-tracking itself is not the challenge in this dataset.
+Rather self-prompting is critical in Alfworld.
+Which indicates that long-range reasoning depends on self-prompting.
+Agent Name M-24B Q-7B Q-14B Q-32B G-27B Average
+Baselines
+Act-only (action) 0.29 0.31 0.66 0.87 0.41 0.51
+ReAct (thought + action) 0.44 0.10 0.75 0.89 0.71 0.58
+Our Methods
+StateAct (state+action) 0.09 0.57 0.75 0.76 0.40 0.51
+StateAct (self-prompt+action) 0.36 0.52 0.77 0.90 0.62 0.63
+StateAct (state+react) 0.21 0.44 0.76 0.77 0.73 0.58
+StateAct (self-prompt+react) 0.53 0.39 0.80 0.91 0.64 0.65
+StateAct (self-prompt+state+action) 0.14 0.60 0.71 0.86 0.51 0.56
+StateAct (self-prompt+state+react) 0.49 0.46 0.78 0.90 0.76 0.68
+Table 8: 135 Test Environments from Alfworld.
+Different Columns represent different models.
+In Bold: Best & 2nd Best Solution per Model.
+Light Green Background: Best Solution per Model.
+Dark Green Background: Best Solution Overall.
+Decoding Strategy: Greedy (temperature=0).
+M=Mistral-Instruct-2501, Q=Qwen2.5; G=Gemma 2-Instruct.
+Appendix M Additional results Webshop
+For Webshop, we present the results for ReAct and StateAct. Similarly to Alfworld, we also present the results of StateAct in different forms, see Table 9.
+We can see that our method again outperform the base-agents.
+For most model we see big jumps in improvement using ‘state’ in the Webshop environment, indicating that the environment benefits from additional structured prediction.
+Interestingly in Webshop the results of Act-only are very strong for many models.
+A hypothesis might be that strong instruction tuning introduces these kind of reasoning techniques into the models directly without the need for explicit prompting from the user.
+Additionally, we can see that thoughts often harm performance on Webshop.
+A hypothesis for this is that Webshop already has a lot of textual content so verbose thoughts can be harmful by confusing the model.
+Agent Name M-24B Q-7B Q-14B Q-32B G-27B Average
+Baselines
+act-only (action) 0.32 0.26 0.16 0.11 0.11 0.19
+react (thought + action) 0.34 0.19 0.22 0.27 0.26 0.26
+Our Methods
+StateAct (state+action) 0.38 0.25 0.37 0.29 0.26 0.31
+StateAct (self-prompt+action) 0.00 0.27 0.36 0.23 0.35 0.24
+StateAct (state+react) 0.23 0.14 0.25 0.26 0.04 0.18
+StateAct (self-prompt+react) 0.24 0.14 0.15 0.25 0.14 0.18
+StateAct (self-prompt+state+action) 0.37 0.26 0.36 0.30 0.15 0.29
+StateAct (self-prompt+state+react) 0.35 0.12 0.33 0.32 0.29 0.28
+Table 9: 100 Test Environments from Webshop.
+Different columns represent different models.
+Average is the average of all models.
+Decoding Strategy: Greedy (temperature=0).
+M=Mistral-Instruct-2501, Q=Qwen2.5; G=Gemma 2-Instruct.
+Appendix N Additional results for Textcraft
+We can see that in Textcraft similar to Alfworld.
+The results improve when thoughts are used.
+Interestingly, for Textcraft self-prompting does not yield the highest result.
+A hypothesis for is because the crafting recipe needs to followed very closesly for successful completion therefore additional reminders are not necessary.
+Agent Name M-24B Q-7B Q-14B Q-32B G-27B Average
+Baselines
+act-only (action) 0.37 0.05 0.26 0.42 0.25 0.27
+react (thought + action) 0.33 0.02 0.31 0.31 0.18 0.23
+Our Methods
+StateAct (state+action) 0.34 0.06 0.09 0.36 0.25 0.22
+StateAct (self-prompt+action) 0.42 0.01 0.25 0.41 0.23 0.26
+StateAct (state+react) 0.45 0.12 0.38 0.42 0.34 0.34
+StateAct (self-prompt+react) 0.41 0.09 0.35 0.37 0.29 0.30
+StateAct (self-prompt+state+action) 0.29 0.01 0.16 0.35 0.23 0.21
+StateAct (self-prompt+state+react) 0.40 0.04 0.37 0.40 0.34 0.31
+Table 10: Performance across 100 Textcraft test environments with different models. Decoding Strategy: Greedy (temperature=0). M=Mistral-Instruct-2501, Q=Qwen2.5; G=Gemma 2-Instruct.
+Appendix O Use of AI
+We used coding assistants in small parts using continue.dev  [^8] https://www.continue.dev/, last accessed March 2025.
+and claude-sonnet-3.5.
+Small use of ChatGPT was used for Latex advise and particular phrasing of parts of the text.